### v1

# Non-perturbative approach for scalar particle production in Higgs-$R^{2}$ inflation

Abstract
We investigate the non-perturbative production of scalar particles within the Higgs-$R^{2}$ inflation model, focusing on a massive spectator field that interacts gravitationally via a non-minimal coupling to the Ricci scalar $R$. By transforming the model to the Einstein frame, where direct couplings between inflationary dynamics and the spectator field emerge, we analyze both perturbative and non-perturbative particle production. We solve the spectator field equation numerically, using the Bogolyubov transformation, to calculate the comoving particle density. Our findings indicate that particle production occurs for light particles with conformal coupling and for masses $m_{\chi}>M/2$ in the minimally coupled scenario.
Inflation, multi-field inflation, scalar fields, particle production
††preprint: APS/123-QED
I Introduction
Cosmic inflation is the cornerstone of modern cosmology, providing a framework to understand the early universe and predict its fundamental characteristics. It describes a period of accelerated expansion driven by a scalar field $\phi$ slow-rolling towards the minimum of its potential $V(\phi)$. This mechanism addresses classical problems of the standard cosmological model and has become a central paradigm in cosmology [1, 2]. During the inflationary epoch, significant particle production occurs, especially as the kinetic energy of the inflaton field approaches its potential energy. This particle production can arise from gravitational effects due to the universe’s expansion and the dynamics of the inflaton field, serving as a classical background for propagating quantum fields [3, 4]. Known as gravitational particle production, this is a fundamental mechanism in quantum field theory in curved spacetime, offering insights into the origins of dark matter and the preheating and reheating epochs following inflation [5]. Several scenarios have been proposed for vacuum particle production during the early universe, particularly during inflation due to the evolution of the inflaton field and the universe’s expansion. Typically, the inflaton field is coupled to matter fields (bosonic, fermionic, or gauge fields) by a coupling constant $g$. At a critical value $\phi=\phi_{*}$ of the inflaton field, resonance begins and particle production is instantaneous [6, 7]. However, matter fields need not be directly coupled with the inflaton or standard model particles. Massive particles can be produced purely through gravitational effects, either via direct couplings with the gravitational field or induced couplings through a conformal transformation. If the matter field’s energy density contribution is negligible, the inflationary dynamics remain unaffected, leading to no back-reaction effects on spacetime geometry. Such fields are called spectator fields, and this mechanism provides a natural explanation for the origin of dark matter [8, 9, 10, 11, 12, 13]. Beyond dark matter origins, gravitational particle production is linked to significant early universe phenomena such as preheating after inflation [14, 15, 16, 17], primordial gravitational waves [18, 19, 20] and primordial magnetogenesis [21, 22, 23, 24]. Additionally, it imprints significant observable effects on the Cosmic Microwave Background (CMB), particularly in the generation of primordial non-Gaussianities [25, 26, 27].
In this work, we study gravitational particle production during and immediately after inflation driven by the SM Higgs boson and the scalaron field in the Higgs-$R^{2}$ inflation model. This model of inflation extends the SM Higgs inflationary model [28] to address the problem of large coupling [29, 30]. Here, cosmic inflation is driven by both the Higgs field and the extra scalar degree of freedom from the quadratic $R$ term, known as the scalaron. Unlike other multi-field inflation models, the Higgs-$R^{2}$ model effectively reduces to a single-field model during inflation, with the Higgs field acting as an auxiliary field while the scalaron drives inflation [31]. At the end of inflation during the reheating phase, both fields and the Ricci scalar oscillate around the potential’s minimum. This oscillatory behaviour enhance the scalar particle production through the non-perturbative mechanisms as tachyonic inestabilities [32, 33].
We propose an extension to the Higgs-$R^{2}$ model by introducing a massive scalar field $\chi$ as a spectator field that it not contributing to the energy density during the inflationary epoch with negligible interactions with the SM but considering a direct coupling to Ricci scalar via a non-minimal coupling term $\xi_{\chi}R\,\chi^{2}$. Initially, there is no explicit coupling between the inflationary sector and the spectator field in the Jordan frame; however, gravitational interactions between the scalaron, the Higgs field and free scalars induce couplings when we change the conformal frame. This process results in gravitational particle production with significant implications for preheating after inflation and background dynamics. Induced coupling in the Einstein frame between free scalars and scalar sector of the action provides an excellent channel of scalaron decay into light scalars. However, in the non-perturbative approach the field mass $m_{\chi}$ does not necessarily have to be light and we can explore the possibility of scalar particle production for spectator field masses $m_{\chi}>M/2$ where $M\simeq 1.3\times 10^{-5}\,M_{p}$ is the scalaron mass. We take into account the oscillatory behavior of background dynamics after inflation as a classical background, which results in particle production due to the time-dependent background inducing a time-dependent frequency in the mode equations for the spectator field $\chi$. To achieve this, we use the Bogoliubov approach [3] to compute the comoving number density $n_{k}(t)$ of particles created by the background dynamics in Higgs-$R^{2}$ inflation by solving the mode equations for the quantized modes $\chi$.
The article is organized as follows: section 2 provides a brief introduction to the Higgs-$R^{2}$ model and the background dynamics during and after inflation for a set of suitable initial conditions. Section 3 analyzes the couplings induced by gravitational effects in the Einstein frame, showing the decay ratio of scalaron into light free scalars. This result indicates that the perturbative approach is kinematically valid only for masses $m_{\chi}<M/2$ during the reheating epoch after inflation when $|R|<M$. In section 4 we derive the mode equation for the spectator field, which is necessary to investigate the behaviour of the time-dependent frequency during and after inflation. We adapt a numerical approach to solve the mode equation for modes that leave $5$ e-folds before the horizon crossing during inflation, which ensures that each mode is within the horizon at the start of the numerical simulation.
Section 4 focuses on massive particle production during preheating. Finally, section 5 discusses the results and conclusions. In the following, we will use natural units in wihch ℏ$\hbar=1$, $c=1$, $\kappa_{\mathrm{B}}=1$ and we use the mass Planck as $8\pi G=M_{p}^{-2}$. We also work in the context of a spatially flat, homogeneous and isotropic universe described by the FLRW metric
$\mathrm{d}s^{2}$ $=$ −$-\mathrm{d}t^{2}+a^{2}(t)(\mathrm{d}x^{2}+\mathrm{d}y^{2}+\mathrm{d}z^{2})$
$=$ $a^{2}(\tau)(-\mathrm{d}\tau^{2}+\mathrm{d}x^{2}+\mathrm{d}y^{2}+\mathrm{d}z^{2})\,,$
where $t$ is the cosmic time, $\tau$ the conformal time and the relation between the two is $\mathrm{d}\tau=\mathrm{d}t/a(t)$ with $a(t)$ is the scale factor.
II Brief review of Higgs-$R^{2}$ inflationary model
In this section, we will give a brief review of the Higgs-$R^{2}$ model in both Einstein and Jordan frames. For further details, we recommend the main references [31, 34, 30, 29] and the references therein. Let us get started the discussion in the Jordan frame; the model is giving by the action
$S_{\mathrm{J}}=\int\mathrm{d}x^{4}\sqrt{-g_{J}}\left[\dfrac{1}{2}(M_{p}^{2}+\xi_{h}h^{2})R_{J}+\dfrac{\xi_{s}}{4}\,R_{J}^{2}\right.\\[5.69046pt] \left.-\dfrac{1}{2}g^{\mu\nu}_{J}\,\partial_{\mu}h\,\partial_{\nu}h-\dfrac{\lambda}{4}\,h^{4}\right]\,,$ (2)
where $h$ is the Higgs field in the unitary gauge, the subindex $J$ denotes variables defined in the Jordan frame, $\xi_{h}$ is the constant coupling between Higgs field and the gravitational sector, $\xi_{s}$ is a mass parameter and $\lambda\sim 10^{-3}$ is the quartic self constant coupling of the Higgs field. This model contains a quadratic term in $R_{J}$ described by the constant $\xi_{s}$, wihch comes from the $R^{2}$ model [35] . Such constant is related to the scalaron mass as
$\xi_{s}=\dfrac{M_{p}^{2}}{3M^{2}}\,,$ (3)
where $M\simeq 1.3\times 10^{-5}M_{p}$ in order to be consistent with the value of the amplitude of primordial perturbations $A_{\zeta}$ [36]. The model (2) can be regarded as a UV extension of Higgs inflation, so it is interesting to make a double interpretation of the gravitational coupling in this conformal frame. In Higgs inflation, the constant coupling $\xi_{h}$ must fulfill the condition $\xi_{h}^{2}/\lambda\sim 10^{10}$ in order to be consistent with CMB observations [28, 37]. This causes that the cutoff scale of the model to be below the natural energy scale (Planck scale) during inflation, making the predictions of the model questionable; however, it has been shown that including the Ricci scalar cuadratic term modifies the value of the energy scale, rasinig it up to the Planck scale. To solve the strong copuling problem present in the scalar sector, the parameter $\xi_{s}$ has to obey the inequality [29]
$\xi_{s}\gtrsim\dfrac{\xi_{h}^{2}}{4\pi}\,.$ (4)
This inequality allow us to make a perturbative treatment of the model. In order to describe the inflationary dynamics of the model, it is more convenient to bring the action (2) into the canonical Einstein-Hilbert form. So that we need to carry out a conformal transformation $\Omega^{2}(x)$ introducing an extra scalar degree of freedom $\phi$ called scalaron, which comes from the quadratic term in $R_{J}$. Thus, to go from Jordan frame to Eintein frame one has to perform a conformal transformation of the space-time metric $g_{\mu\nu}\to\Psi^{2}(x)g_{\mu\nu}$ with $\Psi(x)>0$, where the correct form of it is
$\phi=\alpha^{-1}\log\Psi^{2}\,,$ (5)
where $\alpha=\sqrt{2/3}M_{p}^{-1}$. The conformal transformation (5) removes the quadratic term of the scalar curvature $R_{J}$ introducing the scalaron $\phi$. Then in the Einstein frame the Higgs-$R^{2}$ model of inflation takes the form of a two-sacalar fields model with non-canonical kinetic term
$S_{E}=\int\mathrm{d}^{4}x\,\sqrt{-g}\,\left\{\dfrac{M_{p}^{2}}{2}R-\dfrac{1}{2}g^{\mu\nu}\partial_{\mu}\phi\,\partial_{\nu}\phi\right.\\[5.69046pt] \left.-\dfrac{1}{2}\,e^{-\alpha\phi}g^{\mu\nu}\partial_{h}\,\partial_{\nu}h-V(\phi\,,h)\right\}\,,$ (6)
which can be wirtten in the general action for multi-field inflationary models
$S=\int\mathrm{d}^{4}x\,\sqrt{-g}\left[\dfrac{M_{p}^{2}}{2}R\right.\\[5.69046pt] \left.-\dfrac{1}{2}G_{IJ}\,g^{\mu\nu}\partial_{\mu}\phi^{I}\,\partial_{\nu}\phi^{J}-V(\phi^{I})\right]\,,$ (7)
where in our case $\phi^{I}=(\phi\,,h)$ and $G_{IJ}$ is the non-flat field-space metric given by
$G_{IJ}=\begin{pmatrix}1&0\\ 0&e^{-\alpha\phi}\end{pmatrix}\,.$ (8)
We can see that the non-canonical kinetic term is associated with the geometry of the field space via the metric $G_{IJ}$ [38]. On the other hand, the potential in the Einstein frame has interaction terms between the Higgs field $h$ and the scalaron $\phi$, and takes the form
$V(\phi\,,h)=\dfrac{1}{4}\,e^{-2\alpha\phi}\left[\dfrac{M_{p}^{4}}{\xi_{s}}\left(e^{\alpha\phi}-1-\dfrac{\xi_{h}}{M_{p}^{2}}\,h^{2}\right)^{2}+\lambda\,h^{4}\right]\,.$
Figure 1: Higgs-Starobinsky potential $V(\phi\,,h)/M_{p}^{4}$ in the Einstein frame (9). It is clear the position of the hill at $h=0$ and the position of the two valleys.
The potential $V(\phi^{I})$ depends on both the $\phi$ and $h$, and mixes terms which make more difficult analyze analytically the background trayectories. Nevertheless, during inflation we can make an effective approximation of the model such that inflation is driven by the single scalar field $\phi$ [31], and with the following interesting features
  1. 1.
    $V(\phi^{I})$ shows two valleys and one ridge
  2. 2.
    For $\xi_{h}$, $\xi_{s}$ possitive and enough large, the potential presents an attractor behavior
  3. 3.
    If $\phi\gg M_{p}$ with arbitrary initial conditions, the fields $\phi$ and $h$ will always evolve towards one of the valleys and thus beginning the slow-roll evolution
  4. 4.
    If $\phi<<M_{p}$, the potential is reduced to Higgs potential
    $V(\phi\ll M_{p}\,,h)\simeq V(h)=\dfrac{\lambda_{\mathrm{eff}}}{4}\,h^{4}\,,$ (10)
    where $\lambda_{\mathrm{eff}}$ is the effective quartic self-coupling constant defined by
    $\lambda_{\mathrm{eff}}=\lambda+\dfrac{\xi_{h}^{2}}{\xi_{s}}\,.$ (11)
    The condition $\lambda_{\mathrm{eff}}\lesssim 4\pi$ is neccessary for perturbative preheating
  5. 5.
    If $h\ll M_{p}$, the potential is reduced to Starobinsky potential
    $V(\phi\,,h\ll M_{p})\simeq\dfrac{M_{p}^{4}}{4\xi_{s}}(1-e^{-\alpha\phi})^{2}\,.$ (12)
We can see the positions of the valleys and the hill at $h=0$ in the figure 1. The inflationary dynamics ocurr along one of these valleys, showing a universal attractor behavior, so that inevitably the fields $\phi$ and $h$ will always fall into one of the valleys and then drive slow-roll inflation. To find the valleys, we can perform the analysis in both Einstein [31] and Jordan [39] frames. In the Einstein frame in the limit $\phi\gg M_{p}$ (it is required for enough long inflation) and $h\gg v$, where $v=246\,\mathrm{GeV}$, the valleys are found out by the condition
∂$\dfrac{\partial V}{\partial h}=0\,,$ (13)
which implies that the Higgs field is completely determined by the inflaton field $\phi$
$h^{2}_{\mathrm{min}}(\phi)=\dfrac{M_{p}^{2}\xi_{h}}{\xi_{h}^{2}+\lambda\xi_{s}}\,(e^{\alpha\phi}-1)\,.$ (14)
Because of the potential (9) is symmetric respect to transformation $h\to-h$, the two valleys are localizated by the condition (14). Thus, this shows that the Higgs field is an auxiliary field during inflation, so if we put the result (14) back into the Einstein potential (9), we give the effective potential along the valleys
$V_{\mathrm{eff}}(\phi\,,h_{\mathrm{min}})=\dfrac{M_{p}^{4}}{4(\xi_{s}+\xi_{h}^{2}/\lambda)}\,(1-e^{-\alpha\phi})^{2}$ (15)
Notice that the effective potential reduces to Higgs inflation if $\lambda\xi_{s}\ll\xi_{h}^{2}$ while one get $R^{2}$ inflation for $\lambda\xi_{s}\gg\xi_{h}^{2}$. In order to provide the correct amplitude of primordial scalar perturbations, the amplitude of (15) has to fulfill [40]
$\xi_{s}+\dfrac{\xi_{h}^{2}}{\lambda}=\dfrac{N_{*}^{2}}{72\pi^{2}A_{\zeta}}\simeq 2\times 10^{9}\,,$ (16)
Figure 2: We present the numerical solution of the background dynamics after inflation in cosmic time $t$ for the Higgs and scalaron fields within the valley of the potential. We have numerically solved (17), (18), and (19) using the numerical parameters $\xi_{h}=4000$, $\xi_{s}=2\times 10^{9}$ and $\lambda=0.013$ together with initial conditions $\phi(t_{0})=5.7M_{p}$ and $h(t_{0})=0.096M_{p}$.
where $N_{*}\sim 50-60$ is the number of e-folds before the end of inflation, and $A_{\zeta}\sim 2\times 10^{-9}$ is the amplitude of the primordial power spectrum. For $\lambda=0.013$ and $\xi_{h}\,,\xi_{s}>0$ which satisfy (16), the inflationary dynamics is effectively the same as single-field slow-roll inflation. In this paper we will only analyze the standard parameter region $\xi_{h}\,,\xi_{s}\,,\lambda>0$. Specifically, if we have $\xi_{h}\gg 0$ the isocurvature effects of the model are absent and the bending is whithin the weak mixing case $\eta_{\perp}\to 0$ [31], and the isocurvature and adiabatic modes are not coupled so that they evolve independetly in the standard parameter region of Higgs-$R^{2}$ model. Therefore, the description of the dynamics inside the valleys is an effective single-field inflation.
II.1 Background dynamics
What follows is to analyze the dynamics of the model. We assume that during inflation both scalaron $\phi$ and higgs field $h$ are homogeneous and only depend on cosmic time $t$. In this case the equations of motion can be written as
$\ddot{\phi}+3H\dot{\phi}+\dfrac{\partial V}{\partial\phi}$ $=$ −$-\dfrac{\alpha}{2}\,e^{-\alpha\phi}\dot{h}^{2}\,,$ (17)
$\ddot{h}+3H\dot{h}+e^{\alpha\phi}\,\dfrac{\partial V}{\partial h}$ $=$ $\alpha\dot{\phi}\,\dot{h}\,.$ (18)
The Einstein equations give the Friedman equations for $H(t)=\dot{a}/a$
$3M_{p}^{2}\,H^{2}$ $=$ $\dfrac{1}{2}\dot{\phi}^{2}+\dfrac{1}{2}\,e^{-\alpha\phi}\dot{h}^{2}+V(\phi\,,h)\,,$ (19)
$\dot{H}$ $=$ −$-\dfrac{1}{2M_{p}^{2}}\,\left(\dot{\phi}^{2}+e^{-\alpha\phi}\dot{h}^{2}\right)\,.$ (20)
One has to numerically solve the system of equations (17), (18), (19) and (20) in order to explore the inflationary dynamics of the model in the Einstein frame given some combination values of the parameters $\xi_{h}\,,\xi_{s}\,,\lambda$ and given a suitable initial conditions for $\phi(t_{0})$, $h(t_{0})$. The inflation occurs when $\phi\gg M_{p}$ and the value of the auxiliary field $h$ is given by (14) inside one of the valleys of the potential. In the figure 2 we present the oscillatory regime of both fields after inflation in terms of cosmic time for the initial conditions $\phi_{0}=5.7M_{p}$ and $h_{0}\approx 0.096M_{p}$. However, according to the argument we have presented above, during inflation the isocurvature effects are absent and the dynamics inside the valleys is an effective single-field description. Then the scalaron field should fulfill the single-field equation
$\ddot{\phi}+3H\dot{\phi}+\dfrac{\mathrm{d}V_{\mathrm{eff}}}{\mathrm{d}\phi}\simeq 0\,,$ (21)
and the evolution of the Higgs field is determined by the condition (14). We can assume a single-field slow roll approximation in the inflationary dynamics inside the valleys. This happens when $\ddot{\phi}\ll 3H\dot{\phi}$, thus the field equation for $\phi$ is
$3H\dot{\phi}+\dfrac{\mathrm{d}V_{\mathrm{eff}}}{\mathrm{d}\phi}\simeq 0\,,$ (22)
which is intregrable by a quadrature together with the Friedman equations:
$\phi(t)=\phi_{0}+\alpha^{-1}\log\left(1-\dfrac{t}{t_{\mathrm{end}}}\right)\,,$ (23)
$a(t)=e^{H_{0}t}\,\left(1-\dfrac{t}{t_{\mathrm{end}}}\right)^{3/4}\,,$ (24)
$H(t)=H_{0}\,-\dfrac{3}{2}\,t_{\mathrm{end}}^{-1}\,\left(1-\dfrac{t}{t_{\mathrm{end}}}\right)^{-1}\,,$ (25)
where
$t_{\mathrm{end}}=\dfrac{3}{4}\,H_{0}^{-1}\,e^{\alpha\phi_{0}}\,,\quad H_{0}\simeq\dfrac{M_{p}^{4}}{\sqrt{12}(\xi_{s}+\xi_{h}^{2}/\lambda)}\,,$ (26)
is the time that the effective inflation ends and $H_{0}$ is the scale Hubble parameter during inflation, with $\phi_{0}\sim 5.7M_{p}$ is the value of $\phi$ at the beginning of inflation. It is clear from slow-roll solutions that during inflation we have the usual quasi De Sitter solutions for $\phi$, $a$ and $H$. The effective single-field inflation ends when the slow roll parameter $\epsilon$ is $\epsilon(t_{\mathrm{end}})=1$, which happens at the final values
$\phi(t_{\mathrm{end}})$ $=$ $\alpha^{-1}\log\left(1+\dfrac{2}{\sqrt{3}}\right)\approx 0.92M_{p}\,,$ (27)
$H_{e}$$\approx$ $0.54\,H_{0}\,.$ (28)
III Free scalar fields in the Einstein frame
In this section we study spectator fields present in the Higgs-$R^{2}$ model such as free scalar fields, initial assuming no direct coupling between the matter fields and the inflationary sector in the Jordan frame. In other words, we get started in the Jordan frame and after conformal transformation, the matter sector is gravitationally coupled to the inflatonary sector. In the Jordan frame our model is giving by
$S_{J}=\int\mathrm{d}x^{4}\sqrt{-g_{J}}\left[\dfrac{1}{2}(M_{p}^{2}+\xi_{h}H^{2})R_{J}+\dfrac{\xi_{s}}{4}\,R_{J}^{2}\right.\\[5.69046pt] \left.-\dfrac{1}{2}g^{\mu\nu}_{J}\,\partial_{\mu}h\,\partial_{\nu}h-\dfrac{\lambda}{4}\,h^{4}\right]+S_{\mathrm{matt}}[\psi_{i}]\,,$ (29)
where $\psi_{i}$ are the different matter fields that are present during and after inflation. In this work we will consider a massive spectator scalar field $\chi$ that only interacts gravitationally. The action of such field is giving by
$S_{J}[\chi]=\int\mathrm{d}x^{4}\sqrt{-g_{J}}\left(-\dfrac{1}{2}g^{\mu\nu}_{J}\partial_{\mu}\chi\partial_{\nu}\chi-\dfrac{m_{\chi}^{2}}{2}\chi^{2}\right.\\[5.69046pt] \left.-\dfrac{1}{2}\,\xi_{\chi}\,R_{J}\,\chi^{2}\right)\,.$ (30)
The dimensionless parameter $\xi_{\chi}$ represents the non-minimal coupling to gravity sector and may take any value; however, studies based on electroweak vacuum stability [41] suggest that the value of non-minimal coupling constant should not exceed the value $10$. Moreover, there are two interest special cases: $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$. The value $\xi_{\chi}=0$ is called minimal coupling case and the value $\xi_{\chi}=1/6$ is called conformal coupling. The first case corresponds to a free scalar field theory and the second one ensures the conformal invariance of the model if in addition $m_{\chi}=0$. Although there are no direct coupling between matter fields and the inflatonary sector in the first case, both scalaron $\phi$ and Higgs field $h$ gravitationally interact with every matter field is present during inflation and after inflation, and may occur gravitational particle production process. In the second one, free scalar $\chi$ is coupled to the Ricci scalar and the particle production may also ocurr as long as the matter fields are not conformally invariant, namely, if the scalar fied is massive $m_{\chi}\neq 0$. These kind of models which in the scalar fields interacts only gravitationally is specially interesting since it can explain the particle production process, with the gravitational field responsible for creating $\chi$ particles. This process of particle creation due to gravitational interactions has implications in the reheating process [17, 42, 43] and gravitational dark matter production [44, 45, 46, 12, 11, 9, 47].
What follows is to perform the conformal transformation (5) to the action (30) to change to the Einstein frame. Since $\chi$ is a spectator field, it remains energetically subdominant, $\rho_{\chi}\ll 3H^{2}M_{p}^{2}$, throughout the entire background evolution. Therefore, we can apply the same conformal transformation of the metric (5) as in the previous section. By doing so, we avoid considering back-reaction effects on the space-time geometry and the background dynamics. In the Einstein frame, we obtain the following action
$S_{E}[\tilde{\chi}]=-\dfrac{1}{2}\int\mathrm{d}^{4}x\,\sqrt{-g}\left(g^{\mu\nu}\partial_{\mu}\tilde{\chi}\partial_{\nu}\tilde{\chi}+\xi_{\chi}R\tilde{\chi}^{2}\right.\\[8.5359pt] +\left.e^{-\alpha\phi}m_{\chi}^{2}\tilde{\chi}^{2}\right)+3\alpha\left(\xi_{\chi}-\dfrac{1}{6}\right)\int\mathrm{d}^{4}x\,\sqrt{-g}\left[\tilde{\chi}\,g^{\mu\nu}\partial_{\mu}\tilde{\chi}\partial_{\nu}\phi\right.\\[8.5359pt] \left.+\dfrac{\alpha}{4}\tilde{\chi}^{2}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi\right]\,,$ (31)
where $\chi\to\tilde{\chi}=\Psi^{-1}\chi$ is the re-scaled scalar field and the Ricci scalar in the Einstein frame is given by $R=6(\dot{H}+2H^{2})$. We can perform another conformal transformation which involve the scalar field $\tilde{\chi}$ leading to another Einstein frame in order to remove the non-minimal coupling constant $\xi_{\chi}$. However, we are only interested in the situation where $\chi$ is an spectator field and both scalaron field $\phi$ and Higgs field are responsible of the dynamics during and after inflation. One can see that the conformal transformation (5) induces interaction terms between the scalaron $\phi$, the Higgs field $h$ and the free scalar $\tilde{\chi}$. We can identify the interaction terms from the following action
$S_{\mathrm{int}}=3\alpha\left(\xi_{\chi}-\dfrac{1}{6}\right)\int\mathrm{d}^{4}x\,\sqrt{-g}\left[\tilde{\chi}\,g^{\mu\nu}\partial_{\mu}\tilde{\chi}\partial_{\nu}\phi\right.\\[8.5359pt] \left.+\dfrac{\alpha}{4}\tilde{\chi}^{2}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi\right]\,.$ (32)
This action describes the interaction between the background fields $\phi$ and $h$ with the massive spectator field $\tilde{\chi}$. We can see that even with $\xi_{\chi}=0$ this interaction does not disappear, unless the field $\tilde{\chi}$ is conformal invariant ($m_{\chi}=0\,,\xi_{\chi}=1/6$). One can see this fact from the decay ratio of the scalaron into a pair of light scalar particles during reheating epoch [8], [48]
$\Gamma_{\phi\to\chi\chi}=\left(1-6\xi_{\chi}+\dfrac{2m_{\chi}^{2}}{M^{2}}\right)^{2}\,\dfrac{M^{3}\sqrt{1-4m_{\chi}^{2}/M^{2}}}{192\pi M_{p}^{2}}\,.$
If $\xi_{\chi}=1/6$ and $m_{\chi}\to 0$, the decay ratio vanishes. In addition, if the mass of scalar field $\chi$ is $m_{\chi}>M/2$ where $M$ is scalaron mass term, the scalaron decay into a pair of $\chi$ particles is forbidden. Then, the perturbative production of $\chi$ particles is allowed only in the case of light scalars with mass $m_{\chi}<M/2$. However, it can also occur a scalar particle production by non perturbative effects by time-dependent background dynamics of $\phi$, $h$ and $a$ for a spectator field with mass $m_{\chi}>M/2$. In order to study this process it is necessary to find the equation of motion for the spectator field $\tilde{\chi}$ and track the mode evolution during and after inflation. Varying the action (31) respect to $\tilde{\chi}$ yields the following equation of motion
(□−$(\square-e^{-\alpha\phi}\,m_{\chi}^{2}-\xi_{\chi}\,R)\tilde{\chi}\\[5.69046pt] +\left[\dfrac{\alpha}{2}\,\left(1-6\xi_{\chi}\right)\square\phi-\dfrac{\alpha^{2}}{4}\left(1-6\xi_{\chi}\right)(\partial\phi)^{2}\right]\tilde{\chi}=0\,.$ (34)
Notice that the background dynamics is explicit in the equation of the spectator field $\tilde{\chi}$ due to the conformal transformation. This induces a time-dependent effective mass term, $m_{\mathrm{eff}}^{2}(t),$ for the spectator field $\tilde{\chi}$. Consequently, in the Einstein frame, the model presents an interaction term with the dynamics of the classical background through the Ricci scalar, the scale factor $a$, and the dynamics of the fields $\phi$ and $h$. As a result, studying the solutions is generally highly complex from an analytical perspective; hence a numerical treatment of the background dynamics is necessary for a comprehensive analysis of mode evolution, specially after inflation when both fields $\phi$ and $h$ exhibit a strongly interacting anharmonic oscillatory behavior, as one can see in the figure 2. By numerically determining the background evolution of the scalaron, the Higgs field, and the scale factor, we can compute the Ricci scalar using $R=6(\dot{H}+2H^{2})$, ensuring that the dynamics are well-determined. However, during effective inflation along one of the potential’s valleys, we can approximate the background dynamics using the single-field equation of motion (21).
IV Mode evolution
In this section, we numerically solve the equation of motion (34) along with the background equations (17), (18), (I) to obtain the mode functions $\chi$ for calculating the number density of particles produced, $n_{k}$. The equation of motion for $\tilde{\chi}$ considering a FLRW universe (I) can be written in terms of the cosmic time $t$ as
$\ddot{\tilde{\chi}}+3H\dot{\tilde{\chi}}+\Omega^{2}(t)\tilde{\chi}=0\,,$ (35)
where $\Omega^{2}(t)$ is a background variables dependent and is given by
$\Omega^{2}(t)=e^{-\alpha\phi}m_{\chi}^{2}-a^{-2}\nabla^{2}+\xi_{\chi}R\\[5.69046pt] +3\alpha\left(\dfrac{1}{6}-\xi_{\chi}\right)(\ddot{\phi}+3H\dot{\phi})-\dfrac{3\alpha^{2}}{2}\left(\dfrac{1}{6}-\xi_{\chi}\right)\dot{\phi}^{2}\,.$ (36)
Rescaling the field $\tilde{\chi}\to a^{-3/2}(t)\,\tilde{\chi}(t,\mathbf{r})$ and using the background equations of motion (17), (20) the equation for the scalar field $\tilde{\chi}$ becomes a oscillator-like equation with time-dependent frequency
$\ddot{X}(t,\mathbf{r})+\omega^{2}(t)\,X(t,\mathbf{r})=0\,,$ (37)
where $\omega^{2}(t)$ is given by
$\omega^{2}(t)=\Omega^{2}(t)-\dfrac{9}{4}H^{2}-\dfrac{3}{2}\dot{H}\,.$ (38)
Particle production in the early universe is a process of quantum field theory in curved spacetime, where the gravitational field is not quantized, but the free scalar field $\chi$ is quantized. Therefore, the production of $\chi$ particles is a quantum effect [3, 4, 5]. We proceed by quantizing the spectator field in the Heisenberg picture so that the $\chi$ field and its conjugate momentum $\pi_{\chi}=\dot{\chi}$ are treated as field operators. In Fourier space the rescaled field can be written as
$X(t,\mathbf{r})=\int\dfrac{\mathrm{d}^{3}\mathrm{k}}{(2\pi)^{3}}\left(\hat{a}_{\mathbf{k}}\,X_{k}(t)\,e^{-i\mathbf{k}\cdot\mathbf{r}}+\hat{a}^{\dagger}_{\mathbf{k}}\,X^{*}_{k}(t)\,e^{i\mathbf{k}\cdot\mathbf{r}}\right)\,,$
where $\mathbf{k}$ is the comoving wave number and $\hat{a}_{\mathbf{k}}$, $\hat{a}^{\dagger}_{\mathbf{k}}$ are the creation and annihilation operators respectively, which obey the standard commutation relations at equal times $[a_{\mathbf{k}}\,,a^{\dagger}_{\mathbf{k^{\prime}}}]=(2\pi)^{3}\delta^{(3)}(\mathbf{k}-\mathbf{k}^{\prime})$ and $[a_{\mathbf{k}}\,,a_{\mathbf{k^{\prime}}}]=[a^{\dagger}_{\mathbf{k}}\,,a^{\dagger}_{\mathbf{k^{\prime}}}]=0$. The field equation for Fourier mode $X_{k}(t)$ is given by
$\ddot{X}_{k}(t)+\omega_{k}^{2}(t)X_{k}(t)=0\,,$ (40)
with
$\omega_{k}^{2}(t)=\left(\dfrac{k}{a}\right)^{2}+m_{\mathrm{eff}}^{2}(t)\,,$ (41)
and the following effective mass term
$m_{\mathrm{eff}}^{2}(t)=e^{-\alpha\phi}m_{\chi}^{2}+\xi_{\chi}\,R-\dfrac{9}{4}H^{2}-\dfrac{3}{2}\dot{H}\\[5.69046pt] +3\alpha\left(\xi_{\chi}-\dfrac{1}{6}\right)(V_{\phi}-\alpha M_{p}^{2}\dot{H})\,.$ (42)
The equation for $X_{k}$ describes an oscillator with time-dependent frequency $\omega_{k}^{2}(t)$, hence to track the evolution of the modes $X_{k}$ during and after inflation, it is necessary to analyze $\omega_{k}^{2}(t)$. One can see that during the quasi De Sitter phase during inflation when the effective single field inflation is valid, the effective mass is simplified because of $\dot{H}\to 0$. In this case one can see that the effective mass takes the form
$m_{\mathrm{eff}}^{2}(t\ll t_{\mathrm{end}})\simeq e^{-\alpha\phi}m_{\chi}^{2}+12\left(\xi_{\chi}-\dfrac{3}{16}\right)H^{2}\\[5.69046pt] +3\alpha\left(\xi_{\chi}-\dfrac{1}{6}\right)V_{\phi}\,.$ (43)
We can see that the mass term is exponentially suppressed during inflation when $\phi\gg M_{p}$, making the mass of the spectator field much smaller than the Hubble parameter at the end of inflation, $m_{\chi}\ll H_{\mathrm{end}}$. However, after inflation, when the scalaron takes values $\phi\ll M_{p}$, the free scalar $\tilde{\chi}$ becomes superheavy, $m_{\chi}\gg H_{\mathrm{end}}$, due to the exponential dependence on the scalaron $\phi$. Then at the beginning of inflation, when $t\ll t_{\mathrm{end}}$ the Ricci scalar is $R\simeq 12H^{2}$ and the dominant terms in (41) are $k^{2}/a^{2}$, $H^{2}$ and $12\xi_{\chi}H^{2}$. However, towards the end of inflation, the term $\dot{H}$ cannot be neglected, as after inflation, $\dot{H}$ and the background oscillations dominate the dynamics. The term $V_{\phi}$ regulates the oscillations of the effective mass after inflation, but it is small compared to $H^{2}$ and $\dot{H}$. The contribution of each term depends on the moment at which it is considered. The mass term benefits from exponential growth when $\phi\ll M_{p}$, as we anticipated earlier. Therefore, after inflation, the dominant terms are the mass term and $\dot{H}$. Additionally, if the coupling constant is $\xi_{\chi}<1/6$, the effective mass could be negative during inflation. In such a case, the frequency $\omega_{k}^{2}(t)$ can become negative for super-horizon modes $k\ll aH$, leading to a exponential mode growth $\tilde{\chi}\sim e^{|\omega_{k}|t}$ during inflation due to tachyonic instability. This regime may result in a very efficient amplification of particle production [42, 33].
Figure 3: Time-dependent frequency as a function of cosmic time t for a mass $m_{\chi}=10^{-2}M$, a non-minimal coupling $\xi_{\chi}=0$, and $k\gg aH$ for 5 e-folds before horizon crossing. We set $N_{e}=60$ before the end of inflation as the moment when the mode crosses the horizon.
However, for large enough modes $k$, namely
$k^{2}>k_{*}^{2}=-m_{\mathrm{eff}}^{2}(t)\,,$ (44)
have positive frequency $\omega_{k}^{2}(t)>0$ even if $m_{\mathrm{eff}}^{2}<0$, which corresponds to sub-horizon regime $k\gg aH$. This occurs when the modes are well inside the horizon at the beginning of inflation. In the figure 3 we show the evolution of $\omega_{k}^{2}(t)$ for a non-minimal coupling $\xi=0$. Before $t_{\mathrm{cross}}$, the frequency is dominated by the decay of the physical mode $k/a\sim e^{-Ht}$. Once the mode crosses the horizon $(k\ll aH)$, $\omega_{k}^{2}(t)$ remains almost constant, and its value is negative since the coefficient of the dominant term $H^{2}$ is negative. For $t\gtrsim t_{\mathrm{end}}$, the frequency increases and oscillates after inflation as the background terms oscillate around the minimum of the potential (preheating). One can also observe growth due to the exponential dependence of the scalaron in the mass term toward the end of inflation.
To numerically solve equation (40) along with the background equations, we must specify suitable intial conditions that ensure the modes are within the horizon at early times. In the asymptotic limit $t\to 0$ and according to the behavior of $\omega_{k}^{2}(t)$ at the beginning of inflation, the dominant term is $k/a$. In this limit, the modes are sub-horizon $k\gg aH$, and the modes exhibit Minkowski-like behavior. This motivates choosing the Bunch-Davies vacuum as the initial condition [14]
$\lim_{t\to 0}X_{k}(t)=\dfrac{1}{\sqrt{2\omega_{k}}}\,e^{-i\omega_{k}t}\,.$ (45)
[…]
V Description of particle production
Once we have established the behavior of mode functions $X_{k}$ during and after inflation for some values of spectator mass, the non-minimal coupling $\xi_{\chi}$ and comoving wave number $k$, it is turn to analyze the non-perturbative particle production. The production of $\chi$ particles is due to the time dependence of the background variables, in particular the non-adiabatic behaviour of the frequency $\omega_{k}(t)$ that induces a mix among negative- and positive-frequency [3]. The adiabaticity condition is given by the adiabatic paramater $A_{k}$ defined as
Figure 4: An example of non-adiabatic mode evolution leading the gravitational particle production of $\chi$ particles described by the adiabaticity parameter $|A_{k}(t)|$ for $m_{\chi}=10^{-2}M$ and $\xi_{\chi}=1/6$. At the end of inflation, one see an abruptly increase on $|A_{k}(t)|$ and non-adiabaticity is induced, leading gravitational particle production $n_{k}(t)\neq 0$. At late times, when the stabilization occurs, we recover the adiabaticity evolution.
$A_{k}(t)=\dfrac{\dot{\omega}_{k}}{\omega_{k}^{2}}\,.$ (46)
The adiabatic evolution of $\omega_{k}(t)$ under the WKB approximation of the modes $X_{k}(t)$ guarantees that the positive-frequency mode at early times evolves into positive-frequency modes at late times. That is, for asymptotically early times $t\to 0$, the modes $X_{k}(t)$ evolve as [49]
$\lim_{t\to 0}X_{k}(t)=\dfrac{e^{-i\int^{t}\omega_{k}(t^{\prime})\mathrm{d}t^{\prime}}}{\sqrt{2\omega_{k}(t)}}\,.$ (47)
This behavior usually happens at early times $t_{i}\to 0$ and at sufficiently late times $t_{f}\to+\infty$, but for intermediate times $t_{i}<t_{*}<t_{f}$ during the expansion of the universe the evolution of the modes is no longer adiabatic $|A_{k}|\gg 1$. During inflation, the positive-frequency solution corresponds to the Bunch-Davies vacuum (45), as described in the previous section. This evolution remains valid as long as the parameter $|A_{k}|\ll 1$ at all times. While this condition is met, no particle production occurs, as adiabatic evolution is ensured, and there is no mixing of negative- and positive-frequency modes. From figure 3, we can see that the evolution of $\omega_{k}(t)$ is smooth until the end of inflation. At that point, the condition $|A_{k}|\ll 1$ is strongly violated, leading to an enhancement in particle production and gravitational particle production is enhanced at the time $t_{*}$, which corresponds to the end of inflation. We show the non-adiabaticity evolution in the figure 4 parametrized by the adiabatic parameter $|A_{k}(t)|$. The mixture of positive and negative frequency modes at time $t_{\mathrm{end}}$ can be represented as the sum of the solutions $\exp(\pm i\int^{t}\omega_{k}(t^{\prime})\mathrm{d}t^{\prime})$ using the WKB approximation [14, 5]
$X_{k}(t)=\dfrac{\alpha(t)}{\sqrt{2\omega_{k}(t)}}\,e^{-i\int^{t}\omega_{k}(t^{\prime})\mathrm{d}t^{\prime}}+\dfrac{\beta(t)}{\sqrt{2\omega_{k}(t)}}\,e^{i\int^{t}\omega_{k}(t^{\prime})\mathrm{d}t^{\prime}}\,,$
where $\alpha_{k}(t)$ and $\beta_{k}(t)$ are the time-dependent Bogolyubov coefficients. Comparing with the Bunch-Davies initial condition (47), one see thats $\alpha_{k}(t_{0})=1$ and $\beta_{k}(t_{0})=0$. The comoving particle number density $n_{\chi}(t)$ is related to the coefficient $\beta_{k}$ of the Bogolyubov transformations which relate the vacuum states at the asymptotic limits $t\to 0$ and $t\to\infty$, and coincides with $\beta_{k}(t)$ at late times. We can compute $\beta_{k}(t)$ by solving (40) and evaluating
$|\beta_{k}(t)|^{2}=\dfrac{\omega_{k}}{2}|X_{k}|^{2}+\dfrac{|\dot{X}_{k}|^{2}}{2\omega_{k}}-\dfrac{1}{2}\,.$ (49)
Figure 5: Left: Comoving number density of particles created for $m_{\chi}=10^{-1}M$ and $\xi_{\chi}=1/6$. The graph shows the evolution of the occupation number $n_{k}/k_{\mathrm{end}}^{3}$ as a function of time $t$, where the mode $k$ was chosen to be 5 e-folds before leaving the horizon during inflation and $k_{\mathrm{end}}=a_{\mathrm{end}}H_{\mathrm{end}}$ is the mode at the end of inflation. Right: Spectrum of the gravitationally produced scalar particles for the same numerical example. Modes with $k=aH$ are on the horizon scale at the end of inflation. The mass and the momenta $k$ both are given in units of $M$.
The physical number density $n_{\chi}(t)$ in the late-time limit $t\to\infty$ is given by
$a^{3}(t)\,n_{\chi}(t)=\int_{0}^{\infty}\mathrm{d}\log k\,n_{k}(t)\,,\quad n_{k}(t)=\dfrac{k^{3}|\beta_{k}(t)|^{2}}{2\pi^{2}}\,,$
where $n_{k}(t)$ is the comoving number density of particles with comoving wavenumber $k$.
V.1 Numerical results
What follows is to find the spectrum of $\chi$ particles and its evolution in the cosmic time $t$. Hence we numerically solve (40), constructed the Bogolyubov coeffcient $|\beta_{k}(t)|$ and compute the spectrum $n_{k}$ using (50). In the previous section we described the procedure to numerically evaluate the set of equations to ensure that each mode $k$ is within the horizon at the beginning of the evolution, which allows us to use the Bunch-Davies vacuum (45) as a initial condition of the mode equation for $X_{K}(t)$. The initial conditions for the background are set at time $t_{\mathrm{ini}}$, corresponding to the moment when the modes are within the horizon. For example, for $N_{*}=60$ at the time of crossing, the initial time corresponds to $t_{\mathrm{ini}}\approx 28.712$ in units of $[10^{-5}M_{p}^{-1}]$, and the corresponding mode $k=aH$ will be the value of $aH$ at time $t_{\mathrm{ini}}$. Then, the Bogolyubov coefficient $|\beta_{k}(t)|^{2}$ is calculated using (49) and the comoving number density of gravitational particle production is determined by $n_{k}(t)$.
V.1.1 Conformal coupling $\xi_{\chi}=1/6$
First of all, we focus on the conformal coupling $\xi_{\chi}=1/6$ and consider light masses $m_{\chi}\ll M/2$. As first example, we show the the comoving number density of particles $n_{k}(t)$ as a function of time $t$ and the corresponding spectrum for the case of $m_{\chi}=10^{-1}M$. In the left panel of figure 5, we observe that the number of particles created saturates at the end of inflation, stabilizing and becoming constant at late times $t\to\infty$. This growth in $n_{k}$ is due to the non-adiabaticity of $\omega_{k}$ at the end of inflation, as can be seen in the figure 4. The oscillations of $n_{k}(t)$ after inflation are due to resonances at certain moments between the oscillations of the modes $X_{k}$ and the frequency itself. In the right panel, we observe the corresponding spectrum of $n_{k}(t)$. Several remarkable features can be observed. For IR modes $(k\ll aH)$, we see that the spectrum is blue-tilted, reaching a peak at some intermediate value $k_{\mathrm{peak}}$, and then is suppressed for UV modes $(k\gg aH)$ due to inflation, that is, those that remain within the horizon during inflation. Modes with $k\ll k_{\mathrm{end}}$ leaves the horizon during inflation and re-enter later during the reheating epoch after inflation. Therefore, the dynamics of the $X_{k}(t)$ modes are dictated by the effective mass term (43) with a quasi-De Sitter spacetime as background with $H\simeq\mathrm{constant}$. It is important to note that the UV modes cross the horizon just at the end of inflation, leading to the production of these modes at that instant and their subsequent evolution during the preheating epoch after inflation, as seen in the left panel of figure 5. During this epoch, the production occurs mainly due to the oscillatory nature of the scalaron and the Higgs field, as shown in figure 2, through the term $\dot{H}$ and the Ricci scalar $R$ via the non-minimal coupling $\xi_{\chi}=1/6$. The suppression of UV modes occurs because the $\chi$ scalars become supermassive at the end of inflation due to the exponentially scalaron-dependence mass, causing the effective mass to become substantial. As a result, gravitational particle production becomes less efficient, leading to a significant suppression in the particle density number $n_{k}(t)$.
Figure 6 presents the spectrum of the comoving number of particles produced across the range of light masses considered in this study.
Figure 6: Comoving spectrum of gravitational scalar particle production $n_{k}$ for a range of light masses of the field $\chi$ as a function of rescaled horizon modes $k/k_{\mathrm{end}}$. Masses are given in units of scalaron mass $M$. One can seen that the UV tail is suppressed for light masses $m_{\chi}<M/2$.
The numerical results exhibit a blue-tilted behavior in the infrared (IR) modes, with an exponential suppression in the ultraviolet (UV) tail for light masses $m_{\chi}<M/2$. The mass of the field plays a key role in determining the amplitude of $n_{k}$ at the peak momentum $k_{\mathrm{peak}}$, which systematically decreases as $m_{\chi}$ increases [44]. It is important to emphasize that, in this mass range, particle production following inflation is predominantly driven by the decay of the scalaron into pairs of light $\chi$ particles, as described by expression (33). Furthermore, figure 6 indicates that non-perturbative particle production within this mass range remains minimal, suggesting that such production could lead to an overproduction [8, 50]. For the specific cases of $m_{\chi}=0.2M$ and $m_{\chi}=0.3M$, the production is significantly more moderate. The spectra also exhibit a tendency to flatten in the UV region, implying that gravitational particle production becomes increasingly inefficient for larger masses.
V.1.2 Non-minimal coupling $\xi_{\chi}=0$
Figure 7: Comoving spectrum of gravitational scalar particle production $n_{k}$ for $\xi_{\chi}=0$ as a function of rescaled horizon modes $k/k_{\mathrm{end}}$. Masses are given in units of scalaron mass $M$. One can seen that the spectrum is nearly scale-invariant with a blue-tilted behavior for larger masses.
Figure 8: Comoving number density of particles $n_{k}/k_{\mathrm{end}}^{3}$ for $m_{\chi}=10^{-1}M$ and $\xi_{\chi}=0$ as a function of cosmic time $t$.
For the case of $\xi_{\chi}=0$, we have a different situation. During inflation, the effective mass (43) can be negative for light masses $m_{\chi}\ll H_{\mathrm{end}}$
$m_{\mathrm{eff}}^{2}(t)\approx e^{-\alpha\phi}m_{\chi}^{2}-\dfrac{9}{4}H^{2}-\dfrac{\alpha}{2}V_{\phi}<0\,,$
and once the modes cross the horizon, we obtain $\omega_{k}^{2}(t)<0$, leading to tachyonic instabilities as pointed out in the discussion following equation (43). Therefore, particle production can be due to resonances from oscillations or due to the instabilities caused by $\omega_{k}^{2}(t)<0$. In figure 8, we can observe the behavior of $n_{k}(t)$ normalized to its value at the end of inflation as a function of cosmic time. We can see that for the same mass value $m_{\chi}=10^{-1}M$, particle production is significantly larger than in the case of conformal coupling (see figure 5).
The tachyonic growth of the modes $X_{k}(t)$ outside the horizon is the main driver for the enhanced production of minimally coupled particles compared to the $\xi_{\chi}=1/6$ case. We observe that after inflation, $n_{k}(t)$ exhibits an anharmonic oscillatory behavior with maximum production peaks until it reaches a constant value as $t\to\infty$. However, for masses $m_{\chi}\gg H_{\mathrm{end}}$, the dominant terms in the effective mass during inflation are $e^{-\alpha\phi}m_{\chi}^{2}$ and $(k/a)^{2}$, which implies a positive frequency $\omega_{k}^{2}(t)>0$. In figure 7, we show the spectrum of scalar particles created in the case of minimal coupling and masses $m_{\chi}\gtrsim M/2$. As can be seen, the spectrum generated in the IR region is qualitatively different as the mass increases. We observe an almost scale-invariant spectrum that is slightly blue-tilted for masses $m_{\chi}>M/2$.
VI Discussion and remarks
In this study, we investigated the gravitational and non-perturbative production of free scalar particles, $\chi$, in Higgs-$R^{2}$ inflation. We began with a free spectator field, $\chi$, interacting solely through the gravitational term $\xi_{\chi}R$ in the Jordan frame. By applying a conformal transformation to transition into the Einstein frame, direct couplings between the $\chi$ field and the scalaron $\phi$ emerged. Within the context of inflationary dynamics in the Einstein frame, we derived the equation of motion for the $\chi$ field, taking into account the non-minimal coupling with R. We numerically solved this system of equations, imposing Bunch-Davies conditions on the $\chi$ field 5 e-folds before the modes crossed the horizon, to determine the mode functions of the $\chi$ field. From these solutions, we extracted the Bogolyubov coefficient $\beta_{k}(t)$ to calculate the comoving particle density and visualize their spectrum. We explored two distinct scenarios for the non-minimal coupling $\xi_{\chi}$, considering different mass ranges. For $\xi_{\chi}=1/6$ and light masses, we observed efficient particle production at the end of inflation due to the term $e^{-\alpha\phi}m_{\chi}^{2}$, which becomes dominant at that stage. The resulting power spectrum exhibited a blue-tilted behavior, indicating that particle production was dominated by modes near the Hubble radius $(aH)^{-1}$ at the end of inflation, while shorter wavelength modes were suppressed.
In the case of $\xi_{\chi}=1/6$ and for masses close to $m_{\chi}\sim M/2$, the spectrum tends to flatten out without exhibiting a peak in the UV region. This phenomenon is possibly due to the mass term that depends exponentially on the scalaron $\phi$, causing the effective mass to become superheavy at the end of inflation. In contrast, for $\xi_{\chi}=0$, the spectrum remained nearly scale-invariant for masses $m_{\chi}\gtrsim M/2$, with a blue tilt emerging for larger masses. For light masses and $\xi_{\chi}=0$, we find that the comoving number of particles created is significantly higher compared to the case of conformally coupled particles, which is due to the tachyonic growth of the modes $X_{k}(t)$. We have not attempted a detailed exploration of the parameter space of the Higgs-$R^{2}$ inflation model or the parameters of the scalar field $\chi$. Instead, we have taken the Higgs-$R^{2}$ inflation model, whose observables are consistent with Planck, and considered the possibility of non-perturbative particle production through the direct couplings that arise between the $\chi$ field and the inflationary sector when changing the conformal frame. We found that inflation driven by the Higgs field and the scalaron can produce scalar particles in the cases $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$ and for a specific range of masses. This field could potentially contribute to the abundance of dark matter, for which its current abundance would need to be calculated.
There remains much to be done for future work, such as considering the production of spin-$1/2$ and spin-$1$ dark matter, the production of gauge bosons $W^{\pm}_{\mu}$, $Z_{\mu}$, as well as the implications on primordial non-Gaussianities in the CMB induced by particle production. It would also be interesting to study the isocurvature constraints for light scalar field masses in this model. We leave these intriguing topics for future studies.
VII Acknowledgements
We are very grateful to Dr. Marcos Garcia for his insightful comments and input, which made it possible to write the code for this article. We would also like to acknowledge Dr. Andrew Long for providing an example of numerical code. The work of F.P. has been sponsored by CONACHyT-Mexico through a doctoral scholarship. This work was also partially supported by CONACHyT-Mexico through the support CBF2023-2024-1937.

### v2

@@
-Non-perturbative approach for scalar particle production in Higgs-$R^{2}$ inflation
+
 Abstract
-We investigate the non-perturbative production of scalar particles within the Higgs-$R^{2}$ inflation model, focusing on a massive spectator field that interacts gravitationally via a non-minimal coupling to the Ricci scalar $R$.
-By transforming the model to the Einstein frame, where direct couplings between inflationary dynamics and the spectator field emerge, we analyze both perturbative and non-perturbative particle production.
-We solve the spectator field equation numerically, using the Bogolyubov transformation, to calculate the comoving particle density.
-Our findings indicate that particle production occurs for light particles with conformal coupling and for masses $m_{\chi}>M/2$ in the minimally coupled scenario.
-Inflation, multi-field inflation, scalar fields, particle production
-††preprint: APS/123-QED
-I Introduction
+Gravitational particle production in cosmology is a mechanism through which particles of different natures are produced during the very early universe.
+It is a general mechanism that explains how the universe became populated with the particles of the Standard Model after cosmic inflation and may also account for the origin of dark matter.
+In this work, we study the non-perturbative production of massive scalar particles in the Higgs-$R^{2}$ inflation model, a two-field scalar inflation model within the Einstein frame.
+We consider spectator scalar fields that are conformally or minimally coupled to the gravitational field through the curvature scalar $R$ which in turn is a time-dependent function determined by the fields driving inflationary dynamics.
+We numerically compute the production of these particles using the Bogolyubov transformation method for each scenario, aiming to assess the spectrum of the produced particles.
+For both scenarios, we consider light particles with masses $m_{\chi}\ll H_{\mathrm{end}}$ and large masses that exceed the Hubble scale at the end of inflation $m_{\chi}\gtrsim H_{\mathrm{end}}$.
+keywords:
+Particle production, Multi-field inflation, Higgs inflation, Early universe
+1 Introduction
 Cosmic inflation is the cornerstone of modern cosmology, providing a framework to understand the early universe and predict its fundamental characteristics.
-It describes a period of accelerated expansion driven by a scalar field $\phi$ slow-rolling towards the minimum of its potential $V(\phi)$.
+It describes a period of accelerated expansion driven by a scalar field $\phi$ that slowly rolls towards the minimum of its potential $V(\phi)$.
 This mechanism addresses classical problems of the standard cosmological model and has become a central paradigm in cosmology [1, 2].
-During the inflationary epoch, significant particle production occurs, especially as the kinetic energy of the inflaton field approaches its potential energy.
+During the inflationary epoch, significant particle production occurs, particularly when the kinetic energy of the inflaton field becomes comparable to its potential energy.
 This particle production can arise from gravitational effects due to the universe’s expansion and the dynamics of the inflaton field, serving as a classical background for propagating quantum fields [3, 4].
-Known as gravitational particle production, this is a fundamental mechanism in quantum field theory in curved spacetime, offering insights into the origins of dark matter and the preheating and reheating epochs following inflation [5].
+This phenomenon, known as gravitational particle production, is a fundamental mechanism in quantum field theory in curved spacetime, offering insights into the origins of dark matter and the preheating and reheating epochs following inflation [5].
 Several scenarios have been proposed for vacuum particle production during the early universe, particularly during inflation due to the evolution of the inflaton field and the universe’s expansion.
 Typically, the inflaton field is coupled to matter fields (bosonic, fermionic, or gauge fields) by a coupling constant $g$.
-At a critical value $\phi=\phi_{*}$ of the inflaton field, resonance begins and particle production is instantaneous [6, 7].
+At a critical value $\phi=\phi_{*}$ of the inflaton field, resonance is triggered, leading to instantaneous particle production [6, 7].
 However, matter fields need not be directly coupled with the inflaton or standard model particles.
 Massive particles can be produced purely through gravitational effects, either via direct couplings with the gravitational field or induced couplings through a conformal transformation.
-If the matter field’s energy density contribution is negligible, the inflationary dynamics remain unaffected, leading to no back-reaction effects on spacetime geometry.
+If the contribution of the matter field to the energy density is negligible, the inflationary dynamics remain unaffected, leading to no back-reaction effects on spacetime geometry.
 Such fields are called spectator fields, and this mechanism provides a natural explanation for the origin of dark matter [8, 9, 10, 11, 12, 13].
-Beyond dark matter origins, gravitational particle production is linked to significant early universe phenomena such as preheating after inflation [14, 15, 16, 17], primordial gravitational waves [18, 19, 20] and primordial magnetogenesis [21, 22, 23, 24].
-Additionally, it imprints significant observable effects on the Cosmic Microwave Background (CMB), particularly in the generation of primordial non-Gaussianities [25, 26, 27].
+Beyond its role in the origin of dark matter, gravitational particle production is also linked to key early-universe phenomena such as preheating after inflation [14, 15, 16, 17], primordial gravitational waves [18, 19, 20], and primordial magnetogenesis [21, 22, 23, 24].
+Furthermore, it leaves significant observable imprints on the Cosmic Microwave Background (CMB), particularly in the generation of primordial non-Gaussianities [25, 26, 27].
 In this work, we study gravitational particle production during and immediately after inflation driven by the SM Higgs boson and the scalaron field in the Higgs-$R^{2}$ inflation model.
-This model of inflation extends the SM Higgs inflationary model [28] to address the problem of large coupling [29, 30].
-Here, cosmic inflation is driven by both the Higgs field and the extra scalar degree of freedom from the quadratic $R$ term, known as the scalaron.
+This inflation model extends the SM Higgs inflationary model [28] to address the problem of large coupling [29, 30].
+Here, cosmic inflation is driven by the Higgs field and the extra scalar degree of freedom from the quadratic $R$ term, known as the scalaron.
 Unlike other multi-field inflation models, the Higgs-$R^{2}$ model effectively reduces to a single-field model during inflation, with the Higgs field acting as an auxiliary field while the scalaron drives inflation [31].
-At the end of inflation during the reheating phase, both fields and the Ricci scalar oscillate around the potential’s minimum.
-This oscillatory behaviour enhance the scalar particle production through the non-perturbative mechanisms as tachyonic inestabilities [32, 33].
-We propose an extension to the Higgs-$R^{2}$ model by introducing a massive scalar field $\chi$ as a spectator field that it not contributing to the energy density during the inflationary epoch with negligible interactions with the SM but considering a direct coupling to Ricci scalar via a non-minimal coupling term $\xi_{\chi}R\,\chi^{2}$.
-Initially, there is no explicit coupling between the inflationary sector and the spectator field in the Jordan frame; however, gravitational interactions between the scalaron, the Higgs field and free scalars induce couplings when we change the conformal frame.
+At the end of inflation during the reheating phase, both fields and the Ricci scalar oscillate around the potential minimum.
+This oscillatory behavior enhances scalar particle production through non-perturbative mechanisms such as tachyonic instabilities [32, 33].
+We propose an extension to the Higgs-$R^{2}$ model by introducing a massive scalar field $\chi$ as a spectator field that does not contribute to the energy density during the inflationary epoch, with negligible interactions with the SM but considering a direct coupling to the Ricci scalar via a non-minimal coupling term $\xi_{\chi}R\,\chi^{2}$.
+Initially, the inflationary sector and the spectator field are not explicitly coupled in the Jordan frame; however, gravitational interactions between the scalaron, the Higgs field, and free scalars induce couplings upon transforming to the Einstein frame.
 This process results in gravitational particle production with significant implications for preheating after inflation and background dynamics.
-Induced coupling in the Einstein frame between free scalars and scalar sector of the action provides an excellent channel of scalaron decay into light scalars.
-However, in the non-perturbative approach the field mass $m_{\chi}$ does not necessarily have to be light and we can explore the possibility of scalar particle production for spectator field masses $m_{\chi}>M/2$ where $M\simeq 1.3\times 10^{-5}\,M_{p}$ is the scalaron mass.
+Induced coupling in the Einstein frame between free scalars and the scalar sector of the action provides an excellent channel for scalaron decay into light scalars.
+However, in the non-perturbative approach, the field mass $m_{\chi}$ does not necessarily have to be light, and we can explore the possibility of scalar particle production for spectator field masses $m_{\chi}\gtrsim H_{\mathrm{end}}$, where $H_{\mathrm{end}}$ is the Hubble ratio at the end of inflation.
 We take into account the oscillatory behavior of background dynamics after inflation as a classical background, which results in particle production due to the time-dependent background inducing a time-dependent frequency in the mode equations for the spectator field $\chi$.
-To achieve this, we use the Bogoliubov approach [3] to compute the comoving number density $n_{k}(t)$ of particles created by the background dynamics in Higgs-$R^{2}$ inflation by solving the mode equations for the quantized modes $\chi$.
+To achieve this, we employ the Bogoliubov approach [3] to compute the comoving number density $n_{k}(t)$ of particles generated by the background dynamics in Higgs-$R^{2}$ inflation, solving the mode equations for the quantized modes $\chi$.
 The article is organized as follows: Section 2 provides a brief introduction to the Higgs-$R^{2}$ model and the background dynamics during and after inflation for a set of suitable initial conditions.
-Section 3 analyzes the couplings induced by gravitational effects in the Einstein frame, showing the decay ratio of scalaron into light free scalars.
-This result indicates that the perturbative approach is kinematically valid only for masses $m_{\chi}<M/2$ during the reheating epoch after inflation when $|R|<M$.
-In section 4 we derive the mode equation for the spectator field, which is necessary to investigate the behaviour of the time-dependent frequency during and after inflation.
-We adapt a numerical approach to solve the mode equation for modes that leave $5$ e-folds before the horizon crossing during inflation, which ensures that each mode is within the horizon at the start of the numerical simulation.
-Section 4 focuses on massive particle production during preheating.
+Section 3 analyzes the couplings induced by gravitational effects in the Einstein frame, in Section 4, we derive the mode equation for the spectator field, which is necessary to investigate the behavior of the time-dependent frequency during and after inflation.
+We adopt a numerical approach to solve the mode equation for modes that leave 5 e-folds before the horizon crossing during inflation, ensuring that each mode is within the horizon at the start of the numerical simulation.
+Section 4 examines the production of massive particles during preheating.
 Finally, Section 5 discusses the results and conclusions.
-In the following, we will use natural units in wihch ℏ$\hbar=1$, $c=1$, $\kappa_{\mathrm{B}}=1$ and we use the mass Planck as $8\pi G=M_{p}^{-2}$.
-We also work in the context of a spatially flat, homogeneous and isotropic universe described by the FLRW metric
-$\mathrm{d}s^{2}$ $=$ −$-\mathrm{d}t^{2}+a^{2}(t)(\mathrm{d}x^{2}+\mathrm{d}y^{2}+\mathrm{d}z^{2})$
+In the following, we will use natural units in which $\hbar=1$, $c=1$, $\kappa_{\mathrm{B}}=1$ and we use the mass Planck as $8\pi G=M_{p}^{-2}$.
+We also work in the context of a spatially flat, homogeneous, and isotropic universe described by the FLRW metric
+$\mathrm{d}s^{2}$ $=$ $-\mathrm{d}t^{2}+a^{2}(t)(\mathrm{d}x^{2}+\mathrm{d}y^{2}+\mathrm{d}z^{2})$
 $=$ $a^{2}(\tau)(-\mathrm{d}\tau^{2}+\mathrm{d}x^{2}+\mathrm{d}y^{2}+\mathrm{d}z^{2})\,,$
 where $t$ is the cosmic time, $\tau$ the conformal time and the relation between the two is $\mathrm{d}\tau=\mathrm{d}t/a(t)$ with $a(t)$ is the scale factor.
-II Brief review of Higgs-$R^{2}$ inflationary model
+2 Brief review of Higgs-$R^{2}$ inflationary model
 In this section, we will give a brief review of the Higgs-$R^{2}$ model in both Einstein and Jordan frames.
 For further details, we recommend the main references [31, 34, 30, 29] and the references therein.
-Let us get started the discussion in the Jordan frame; the model is giving by the action
-$S_{\mathrm{J}}=\int\mathrm{d}x^{4}\sqrt{-g_{J}}\left[\dfrac{1}{2}(M_{p}^{2}+\xi_{h}h^{2})R_{J}+\dfrac{\xi_{s}}{4}\,R_{J}^{2}\right.\\[5.69046pt] \left.-\dfrac{1}{2}g^{\mu\nu}_{J}\,\partial_{\mu}h\,\partial_{\nu}h-\dfrac{\lambda}{4}\,h^{4}\right]\,,$ (2)
-where $h$ is the Higgs field in the unitary gauge, the subindex $J$ denotes variables defined in the Jordan frame, $\xi_{h}$ is the constant coupling between Higgs field and the gravitational sector, $\xi_{s}$ is a mass parameter and $\lambda\sim 10^{-3}$ is the quartic self constant coupling of the Higgs field.
-This model contains a quadratic term in $R_{J}$ described by the constant $\xi_{s}$, wihch comes from the $R^{2}$ model [35] .
-Such constant is related to the scalaron mass as
-$\xi_{s}=\dfrac{M_{p}^{2}}{3M^{2}}\,,$ (3)
-where $M\simeq 1.3\times 10^{-5}M_{p}$ in order to be consistent with the value of the amplitude of primordial perturbations $A_{\zeta}$ [36].
+Let us get started with the discussion in the Jordan frame; the model is given by the action
+$S_{\mathrm{J}}=\int\mathrm{d}x^{4}\sqrt{-g_{J}}\left[\dfrac{1}{2}(M_{p}^{2}+\xi_{h}h^{2})R_{J}+\dfrac{\xi_{s}}{4}\,R_{J}^{2}-\dfrac{1}{2}g^{\mu\nu}_{J}\,\partial_{\mu}h\,\partial_{\nu}h-\dfrac{\lambda}{4}\,h^{4}\right]\,,$ (2)
+where $h$ is the Higgs field in the unitary gauge, the subindex $J$ denotes variables defined in the Jordan frame, $\xi_{h}$ is the constant coupling between the Higgs field and the gravitational sector, $\xi_{s}$ is a mass parameter and $\lambda$ is the quartic self constant coupling of the Higgs field.
+This model contains a quadratic term in $R_{J}$ described by the constant $\xi_{s}$, which comes from the $R^{2}$ model [35].
+Such a constant is related to the scalaron mass as $\xi_{s}=M_{p}^{2}/3M^{2}$.
 The model (2) can be regarded as a UV extension of Higgs inflation, so it is interesting to make a double interpretation of the gravitational coupling in this conformal frame.
-In Higgs inflation, the constant coupling $\xi_{h}$ must fulfill the condition $\xi_{h}^{2}/\lambda\sim 10^{10}$ in order to be consistent with CMB observations [28, 37].
-This causes that the cutoff scale of the model to be below the natural energy scale (Planck scale) during inflation, making the predictions of the model questionable; however, it has been shown that including the Ricci scalar cuadratic term modifies the value of the energy scale, rasinig it up to the Planck scale.
-To solve the strong copuling problem present in the scalar sector, the parameter $\xi_{s}$ has to obey the inequality [29]
+In Higgs inflation, the constant coupling $\xi_{h}$ must fulfill the condition $\xi_{h}^{2}/\lambda\sim 10^{10}$ to be consistent with CMB observations [28, 36].
+This causes the cutoff scale of the model to be below the natural energy scale (Planck scale) during inflation, making the predictions of the model questionable; however, it has been shown that including the Ricci scalar quadratic term modifies the value of the energy scale, raising it to the Planck scale.
+To solve the strong coupling problem present in the scalar sector, the parameter $\xi_{s}$ has to obey the relation [29]
 $\xi_{s}\gtrsim\dfrac{\xi_{h}^{2}}{4\pi}\,.$ (3)
-This inequality allow us to make a perturbative treatment of the model.
-In order to describe the inflationary dynamics of the model, it is more convenient to bring the action (2) into the canonical Einstein-Hilbert form.
-So that we need to carry out a conformal transformation $\Omega^{2}(x)$ introducing an extra scalar degree of freedom $\phi$ called scalaron, which comes from the quadratic term in $R_{J}$.
-Thus, to go from Jordan frame to Eintein frame one has to perform a conformal transformation of the space-time metric $g_{\mu\nu}\to\Psi^{2}(x)g_{\mu\nu}$ with $\Psi(x)>0$, where the correct form of it is
+This expression allows us to make a perturbative treatment of the model.
+To describe the inflationary dynamics of the model, it is more convenient to bring the action (2) into the canonical Einstein-Hilbert form.
+So we need to carry out a conformal transformation $\Omega^{2}(x)$ introducing an extra scalar degree of freedom $\phi$ called scalaron, which comes from the quadratic term in $R_{J}$.
+Thus, to go from the Jordan frame to Einstein frame one has to perform a conformal transformation of the space-time metric $g_{\mu\nu}\to\Psi^{2}(x)g_{\mu\nu}$ with $\Psi(x)>0$, where the correct form of it is
 $\phi=\alpha^{-1}\log\Psi^{2}\,,$ (4)
@@
 The conformal transformation (4) removes the quadratic term of the scalar curvature $R_{J}$ introducing the scalaron $\phi$.
-Then in the Einstein frame the Higgs-$R^{2}$ model of inflation takes the form of a two-sacalar fields model with non-canonical kinetic term
-$S_{E}=\int\mathrm{d}^{4}x\,\sqrt{-g}\,\left\{\dfrac{M_{p}^{2}}{2}R-\dfrac{1}{2}g^{\mu\nu}\partial_{\mu}\phi\,\partial_{\nu}\phi\right.\\[5.69046pt] \left.-\dfrac{1}{2}\,e^{-\alpha\phi}g^{\mu\nu}\partial_{h}\,\partial_{\nu}h-V(\phi\,,h)\right\}\,,$ (6)
-which can be wirtten in the general action for multi-field inflationary models
-$S=\int\mathrm{d}^{4}x\,\sqrt{-g}\left[\dfrac{M_{p}^{2}}{2}R\right.\\[5.69046pt] \left.-\dfrac{1}{2}G_{IJ}\,g^{\mu\nu}\partial_{\mu}\phi^{I}\,\partial_{\nu}\phi^{J}-V(\phi^{I})\right]\,,$ (7)
+Then in the Einstein frame, the Higgs-$R^{2}$ model of inflation takes the form of a two-scalar fields model with a non-canonical kinetic term
+$S_{E}=\int\mathrm{d}^{4}x\,\sqrt{-g}\,\left\{\dfrac{M_{p}^{2}}{2}R-\dfrac{1}{2}g^{\mu\nu}\partial_{\mu}\phi\,\partial_{\nu}\phi-\dfrac{1}{2}\,e^{-\alpha\phi}g^{\mu\nu}\partial_{h}\,\partial_{\nu}h-V(\phi\,,h)\right\}\,,$ (5)
+which can be written in the general action for multi-field inflationary models
+Figure 1: Scalar potential in the Einstein frame of the Higgs-$R^{2}$ model seen from different angles.
+The curve $\phi=\mathbf{cte}$ has the shape of the potential of the $R^{2}$ inflation model, while $h=\mathrm{cte}$ has the shape of the quartic Higgs potential of the Standard Model.
+We have restricted ourselves to the standard parameter space $\xi_{h}=4000$, $\xi_{s}=2\times 10^{9}$ and $\lambda=0.013$.
+$S=\int\mathrm{d}^{4}x\,\sqrt{-g}\left[\dfrac{M_{p}^{2}}{2}R-\dfrac{1}{2}G_{IJ}\,g^{\mu\nu}\partial_{\mu}\phi^{I}\,\partial_{\nu}\phi^{J}-V(\phi^{I})\right]\,,$ (6)
 where in our case $\phi^{I}=(\phi\,,h)$ and $G_{IJ}$ is the non-flat field-space metric given by
@@
 We can see that the non-canonical kinetic term is associated with the geometry of the field space via the metric $G_{IJ}$ [37].
-On the other hand, the potential in the Einstein frame has interaction terms between the Higgs field $h$ and the scalaron $\phi$, and takes the form
+On the other hand, the potential in the Einstein frame has interaction terms between the Higgs field $h$ and the scalaron $\phi$ and takes the form
 $V(\phi\,,h)=\dfrac{1}{4}\,e^{-2\alpha\phi}\left[\dfrac{M_{p}^{4}}{\xi_{s}}\left(e^{\alpha\phi}-1-\dfrac{\xi_{h}}{M_{p}^{2}}\,h^{2}\right)^{2}+\lambda\,h^{4}\right]\,.$
-Figure 1: Higgs-Starobinsky potential $V(\phi\,,h)/M_{p}^{4}$ in the Einstein frame (9). It is clear the position of the hill at $h=0$ and the position of the two valleys.
-The potential $V(\phi^{I})$ depends on both the $\phi$ and $h$, and mixes terms which make more difficult analyze analytically the background trayectories.
-Nevertheless, during inflation we can make an effective approximation of the model such that inflation is driven by the single scalar field $\phi$ [31], and with the following interesting features
+The potential $V(\phi^{I})$ depends on both the $\phi$ and $h$ and mixes terms which make it more difficult to analyze analytically the background trajectory.
+Nevertheless, during inflation, we can make an effective approximation of the model such that inflation is driven by the single scalar field $\phi$ [31] and with the following interesting features
   1. 1.
-    $V(\phi^{I})$ shows two valleys and one ridge
+    $V(\phi^{I})$ shows two valleys and one ridge.
   2. 2.
-    For $\xi_{h}$, $\xi_{s}$ possitive and enough large, the potential presents an attractor behavior
+    For $\xi_{h}$, $\xi_{s}$ positive and enough large, the potential presents an attractor behavior.
   3. 3.
-    If $\phi\gg M_{p}$ with arbitrary initial conditions, the fields $\phi$ and $h$ will always evolve towards one of the valleys and thus beginning the slow-roll evolution
+    If $\phi\gg M_{p}$ with arbitrary initial conditions, the fields $\phi$ and $h$ will always evolve towards one of the valleys thus beginning the slow-roll evolution.
   4. 4.
-    If $\phi<<M_{p}$, the potential is reduced to Higgs potential
+    If $\phi\ll M_{p}$, the potential is reduced to Higgs potential
     $V(\phi\ll M_{p}\,,h)\simeq V(h)=\dfrac{\lambda_{\mathrm{eff}}}{4}\,h^{4}\,,$ (9)
@@
     $\lambda_{\mathrm{eff}}=\lambda+\dfrac{\xi_{h}^{2}}{\xi_{s}}\,.$ (10)
-    The condition $\lambda_{\mathrm{eff}}\lesssim 4\pi$ is neccessary for perturbative preheating
+    The condition $\lambda_{\mathrm{eff}}\lesssim 4\pi$ is necessary for perturbative preheating
   5. 5.
@@
     $V(\phi\,,h\ll M_{p})\simeq\dfrac{M_{p}^{4}}{4\xi_{s}}(1-e^{-\alpha\phi})^{2}\,.$ (11)
-We can see the positions of the valleys and the hill at $h=0$ in the figure 1.
-The inflationary dynamics ocurr along one of these valleys, showing a universal attractor behavior, so that inevitably the fields $\phi$ and $h$ will always fall into one of the valleys and then drive slow-roll inflation.
+We can see the positions of the valleys and the hill at $h=0$ in figure 1.
+The inflationary dynamics occur along one of these valleys, showing a universal attractor behavior, so that inevitably the fields $\phi$ and $h$ will always fall into one of the valleys and then drive slow-roll inflation.
 To find the valleys, we can perform the analysis in both Einstein [31] and Jordan [38] frames.
 In the Einstein frame in the limit $\phi\gg M_{p}$ (it is required for enough long inflation) and $h\gg v$, where $v=246\,\mathrm{GeV}$, the valleys are found out by the condition
-∂$\dfrac{\partial V}{\partial h}=0\,,$ (13)
+$\dfrac{\partial V}{\partial h}=0\,,$ (12)
 which implies that the Higgs field is completely determined by the inflaton field $\phi$
 $h^{2}_{\mathrm{min}}(\phi)=\dfrac{M_{p}^{2}\xi_{h}}{\xi_{h}^{2}+\lambda\xi_{s}}\,(e^{\alpha\phi}-1)\,.$ (13)
-Because of the potential (9) is symmetric respect to transformation $h\to-h$, the two valleys are localizated by the condition (14).
+Because the potential (8) is symmetric with respect to transformation $h\to-h$, the two valleys are localized by the condition (13).
 Thus, this shows that the Higgs field is an auxiliary field during inflation, so if we put the result (13) back into the Einstein potential (8), we give the effective potential along the valleys
 $V_{\mathrm{eff}}(\phi\,,h_{\mathrm{min}})=\dfrac{M_{p}^{4}}{4(\xi_{s}+\xi_{h}^{2}/\lambda)}\,(1-e^{-\alpha\phi})^{2}$ (14)
-Notice that the effective potential reduces to Higgs inflation if $\lambda\xi_{s}\ll\xi_{h}^{2}$ while one get $R^{2}$ inflation for $\lambda\xi_{s}\gg\xi_{h}^{2}$.
-In order to provide the correct amplitude of primordial scalar perturbations, the amplitude of (15) has to fulfill [40]
+Notice that the effective potential reduces to Higgs inflation if $\lambda\xi_{s}\ll\xi_{h}^{2}$ while one gets $R^{2}$ inflation for $\lambda\xi_{s}\gg\xi_{h}^{2}$.
+To provide the correct amplitude of primordial scalar perturbations, the amplitude of (14) has to fulfill [39]
 $\xi_{s}+\dfrac{\xi_{h}^{2}}{\lambda}=\dfrac{N_{*}^{2}}{72\pi^{2}A_{\zeta}}\simeq 2\times 10^{9}\,,$ (15)
-Figure 2: We present the numerical solution of the background dynamics after inflation in cosmic time $t$ for the Higgs and scalaron fields within the valley of the potential.
-We have numerically solved (17), (18), and (19) using the numerical parameters $\xi_{h}=4000$, $\xi_{s}=2\times 10^{9}$ and $\lambda=0.013$ together with initial conditions $\phi(t_{0})=5.7M_{p}$ and $h(t_{0})=0.096M_{p}$.
 where $N_{*}\sim 50-60$ is the number of e-folds before the end of inflation, and $A_{\zeta}\sim 2\times 10^{-9}$ is the amplitude of the primordial power spectrum.
 For $\lambda=0.013$ and $\xi_{h}\,,\xi_{s}>0$ which satisfy (15), the inflationary dynamics is effectively the same as single-field slow-roll inflation.
-In this paper we will only analyze the standard parameter region $\xi_{h}\,,\xi_{s}\,,\lambda>0$.
-Specifically, if we have $\xi_{h}\gg 0$ the isocurvature effects of the model are absent and the bending is whithin the weak mixing case $\eta_{\perp}\to 0$ [31], and the isocurvature and adiabatic modes are not coupled so that they evolve independetly in the standard parameter region of Higgs-$R^{2}$ model.
+In this paper, we will only analyze the standard parameter region $\xi_{h}\,,\xi_{s}\,,\lambda>0$.
+Specifically, if we have $\xi_{h}\gg 0$ the isocurvature effects of the model are absent the bending is within the weak mixing case $\eta_{\perp}\to 0$ [31], and the isocurvature and adiabatic modes are not coupled so that they evolve independently in the standard parameter region of Higgs-$R^{2}$ model.
 Therefore, the description of the dynamics inside the valleys is an effective single-field inflation.
-II.1 Background dynamics
+2.1 Background dynamics
 What follows is to analyze the dynamics of the model.
 We assume that during inflation both scalaron $\phi$ and Higgs field $h$ are homogeneous and only depend on cosmic time $t$.
-In this case the equations of motion can be written as
-$\ddot{\phi}+3H\dot{\phi}+\dfrac{\partial V}{\partial\phi}$ $=$ −$-\dfrac{\alpha}{2}\,e^{-\alpha\phi}\dot{h}^{2}\,,$ (17)
+In this case, the equations of motion can be written as
+$\ddot{\phi}+3H\dot{\phi}+\dfrac{\partial V}{\partial\phi}$ $=$ $-\dfrac{\alpha}{2}\,e^{-\alpha\phi}\dot{h}^{2}\,,$ (16)
 $\ddot{h}+3H\dot{h}+e^{\alpha\phi}\,\dfrac{\partial V}{\partial h}$ $=$ $\alpha\dot{\phi}\,\dot{h}\,.$ (17)
@@
 $3M_{p}^{2}\,H^{2}$ $=$ $\dfrac{1}{2}\dot{\phi}^{2}+\dfrac{1}{2}\,e^{-\alpha\phi}\dot{h}^{2}+V(\phi\,,h)\,,$ (18)
-$\dot{H}$ $=$ −$-\dfrac{1}{2M_{p}^{2}}\,\left(\dot{\phi}^{2}+e^{-\alpha\phi}\dot{h}^{2}\right)\,.$ (20)
-One has to numerically solve the system of equations (17), (18), (19) and (20) in order to explore the inflationary dynamics of the model in the Einstein frame given some combination values of the parameters $\xi_{h}\,,\xi_{s}\,,\lambda$ and given a suitable initial conditions for $\phi(t_{0})$, $h(t_{0})$.
+$\dot{H}$ $=$ $-\dfrac{1}{2M_{p}^{2}}\,\left(\dot{\phi}^{2}+e^{-\alpha\phi}\dot{h}^{2}\right)\,.$ (19)
+Figure 2: We present the numerical solution of the background dynamics after inflation in cosmic time $t$ for the Higgs and scalaron fields within the valley of the potential.
+We have numerically solved (16), (17), and (18) using the numerical parameters $\xi_{h}=4000$, $\xi_{s}=2\times 10^{9}$ and $\lambda=0.013$ together with initial conditions $\phi(t_{0})=5.7M_{p}$ and $h(t_{0})=0.096M_{p}$.
+One has to numerically solve the system of equations (16), (17), (18) and (19) to explore the inflationary dynamics of the model in the Einstein frame given some combination values of the parameters $\xi_{h}\,,\xi_{s}\,,\lambda$ and given suitable initial conditions for $\phi(t_{0})$, $h(t_{0})$.
 The inflation occurs when $\phi\gg M_{p}$ and the value of the auxiliary field $h$ is given by (13) inside one of the valleys of the potential.
 In the figure 2 we present the oscillatory regime of both fields after inflation in terms of cosmic time for the initial conditions $\phi_{0}=5.7M_{p}$ and $h_{0}\approx 0.096M_{p}$.
-However, according to the argument we have presented above, during inflation the isocurvature effects are absent and the dynamics inside the valleys is an effective single-field description.
+However, according to the argument we have presented above, during inflation the isocurvature effects are absent and the dynamics inside the valleys are an effective single-field description.
 Then the scalaron field should fulfill the single-field equation
@@
 $3H\dot{\phi}+\dfrac{\mathrm{d}V_{\mathrm{eff}}}{\mathrm{d}\phi}\simeq 0\,,$ (21)
-which is intregrable by a quadrature together with the Friedman equations:
+which can be integrated by quadrature together with the Friedman equations
 $\phi(t)=\phi_{0}+\alpha^{-1}\log\left(1-\dfrac{t}{t_{\mathrm{end}}}\right)\,,$ (22)
@@
 is the time that the effective inflation ends and $H_{0}$ is the scale Hubble parameter during inflation, with $\phi_{0}\sim 5.7M_{p}$ is the value of $\phi$ at the beginning of inflation.
-It is clear from slow-roll solutions that during inflation we have the usual quasi De Sitter solutions for $\phi$, $a$ and $H$.
+It is clear from slow-roll solutions that during inflation we have the usual quasi-De Sitter solutions for $\phi$, $a$, and $H$.
 The effective single-field inflation ends when the slow roll parameter $\epsilon$ is $\epsilon(t_{\mathrm{end}})=1$, which happens at the final values
@@
 $H_{e}$$\approx$ $0.54\,H_{0}\,.$ (27)
-III Free scalar fields in the Einstein frame
-In this section we study spectator fields present in the Higgs-$R^{2}$ model such as free scalar fields, initial assuming no direct coupling between the matter fields and the inflationary sector in the Jordan frame.
-In other words, we get started in the Jordan frame and after conformal transformation, the matter sector is gravitationally coupled to the inflatonary sector.
-In the Jordan frame our model is giving by
-$S_{J}=\int\mathrm{d}x^{4}\sqrt{-g_{J}}\left[\dfrac{1}{2}(M_{p}^{2}+\xi_{h}H^{2})R_{J}+\dfrac{\xi_{s}}{4}\,R_{J}^{2}\right.\\[5.69046pt] \left.-\dfrac{1}{2}g^{\mu\nu}_{J}\,\partial_{\mu}h\,\partial_{\nu}h-\dfrac{\lambda}{4}\,h^{4}\right]+S_{\mathrm{matt}}[\psi_{i}]\,,$ (29)
+3 Scalar fields in the Einstein frame
+In this section, we study spectator fields present in the Higgs-$R^{2}$ model such as free scalar fields, initially assuming no direct coupling between the matter fields and the inflationary sector in the Jordan frame.
+In other words, we get started in the Jordan frame and after conformal transformation, the matter sector is gravitationally coupled to the inflationary sector.
+In the Jordan frame our model is given by
+$S_{J}=\int\mathrm{d}x^{4}\sqrt{-g_{J}}\left[\dfrac{1}{2}(M_{p}^{2}+\xi_{h}H^{2})R_{J}+\dfrac{\xi_{s}}{4}\,R_{J}^{2}-\dfrac{1}{2}g^{\mu\nu}_{J}\,\partial_{\mu}h\,\partial_{\nu}h-\dfrac{\lambda}{4}\,h^{4}\right]+S_{\mathrm{matt}}[\psi_{i}]\,$
 where $\psi_{i}$ are the different matter fields that are present during and after inflation.
-In this work we will consider a massive spectator scalar field $\chi$ that only interacts gravitationally.
-The action of such field is giving by
-$S_{J}[\chi]=\int\mathrm{d}x^{4}\sqrt{-g_{J}}\left(-\dfrac{1}{2}g^{\mu\nu}_{J}\partial_{\mu}\chi\partial_{\nu}\chi-\dfrac{m_{\chi}^{2}}{2}\chi^{2}\right.\\[5.69046pt] \left.-\dfrac{1}{2}\,\xi_{\chi}\,R_{J}\,\chi^{2}\right)\,.$ (30)
-The dimensionless parameter $\xi_{\chi}$ represents the non-minimal coupling to gravity sector and may take any value; however, studies based on electroweak vacuum stability [41] suggest that the value of non-minimal coupling constant should not exceed the value $10$.
-Moreover, there are two interest special cases: $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$.
-The value $\xi_{\chi}=0$ is called minimal coupling case and the value $\xi_{\chi}=1/6$ is called conformal coupling.
-The first case corresponds to a free scalar field theory and the second one ensures the conformal invariance of the model if in addition $m_{\chi}=0$.
-Although there are no direct coupling between matter fields and the inflatonary sector in the first case, both scalaron $\phi$ and Higgs field $h$ gravitationally interact with every matter field is present during inflation and after inflation, and may occur gravitational particle production process.
-In the second one, free scalar $\chi$ is coupled to the Ricci scalar and the particle production may also ocurr as long as the matter fields are not conformally invariant, namely, if the scalar fied is massive $m_{\chi}\neq 0$.
-These kind of models which in the scalar fields interacts only gravitationally is specially interesting since it can explain the particle production process, with the gravitational field responsible for creating $\chi$ particles.
-This process of particle creation due to gravitational interactions has implications in the reheating process [17, 42, 43] and gravitational dark matter production [44, 45, 46, 12, 11, 9, 47].
+In this work, we will consider a massive spectator scalar field $\chi$ that only interacts gravitationally.
+The action of such a field is given by
+$S_{J}[\chi]=\int\mathrm{d}x^{4}\sqrt{-g_{J}}\left(-\dfrac{1}{2}g^{\mu\nu}_{J}\partial_{\mu}\chi\partial_{\nu}\chi-\dfrac{m_{\chi}^{2}}{2}\chi^{2}\dfrac{1}{2}\,\xi_{\chi}\,R_{J}\,\chi^{2}\right)\,.$ (29)
+The dimensionless parameter $\xi_{\chi}$ represents the non-minimal coupling to the gravity sector and may take any value; however, studies based on electroweak vacuum stability [40] suggest that the value of non-minimal coupling constant should not exceed $10$.
+Moreover there are two interesting special cases: $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$.
+The value $\xi_{\chi}=0$ is called the case of minimal coupling.
+while $\xi_{\chi}=1/6$ corresponds to conformal coupling.
+The first case corresponds to a free scalar field theory and the second one ensures the conformal invariance of the model if, in addition, $m_{\chi}=0$.
+Although there is no direct coupling between matter fields and the inflationary sector in the first case, both scalaron $\phi$ and Higgs field $h$ gravitationally interact with every matter field present during inflation and after inflation, and a gravitational particle production process may occur.
+In the second one, free scalar $\chi$ is coupled to the Ricci scalar and the particle production may also occur as long as the matter fields are not conformally invariant, namely, if the scalar field is massive $m_{\chi}\neq 0$.
+These kinds of models, in which the scalar fields interact only gravitationally, are especially interesting.
+They can explain the particle production process, with the gravitational field responsible for creating $\chi$ particles.
+This process of particle creation due to gravitational interactions has implications for the reheating process [17, 41, 42].
+It also plays a role in gravitational dark matter production [43, 44, 45, 12, 11, 9, 46].
 What follows is to perform the conformal transformation (4) to the action (29) to change to the Einstein frame.
@@
 where $\chi\to\tilde{\chi}=\Psi^{-1}\chi$ is the re-scaled scalar field and the Ricci scalar in the Einstein frame is given by $R=6(\dot{H}+2H^{2})$.
-We can perform another conformal transformation which involve the scalar field $\tilde{\chi}$ leading to another Einstein frame in order to remove the non-minimal coupling constant $\xi_{\chi}$.
-However, we are only interested in the situation where $\chi$ is an spectator field and both scalaron field $\phi$ and Higgs field are responsible of the dynamics during and after inflation.
-One can see that the conformal transformation (5) induces interaction terms between the scalaron $\phi$, the Higgs field $h$ and the free scalar $\tilde{\chi}$.
+We can perform another conformal transformation which involves the scalar field $\tilde{\chi}$ leading to another Einstein frame to remove the non-minimal coupling constant $\xi_{\chi}$.
+However, we are only interested in the situation where $\chi$ is a spectator field and both scalaron field $\phi$ and Higgs field are responsible for the dynamics during and after inflation.
+One can see that the conformal transformation (4) induces interaction terms between the scalaron $\phi$, the Higgs field $h$, and the free scalar $\tilde{\chi}$.
 We can identify the interaction terms from the following action
-$S_{\mathrm{int}}=3\alpha\left(\xi_{\chi}-\dfrac{1}{6}\right)\int\mathrm{d}^{4}x\,\sqrt{-g}\left[\tilde{\chi}\,g^{\mu\nu}\partial_{\mu}\tilde{\chi}\partial_{\nu}\phi\right.\\[8.5359pt] \left.+\dfrac{\alpha}{4}\tilde{\chi}^{2}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi\right]\,.$ (32)
+$S_{\mathrm{int}}=3\alpha\left(\xi_{\chi}-\dfrac{1}{6}\right)\int\mathrm{d}^{4}x\,\sqrt{-g}\left[\tilde{\chi}\,g^{\mu\nu}\partial_{\mu}\tilde{\chi}\partial_{\nu}\phi+\dfrac{\alpha}{4}\tilde{\chi}^{2}g^{\mu\nu}\partial_{\mu}\phi\,\partial_{\nu}\phi\right]\,.$ (31)
 This action describes the interaction between the background fields $\phi$ and $h$ with the massive spectator field $\tilde{\chi}$.
 We can see that even with $\xi_{\chi}=0$ this interaction does not disappear, unless the field $\tilde{\chi}$ is conformal invariant ($m_{\chi}=0\,,\xi_{\chi}=1/6$).
-One can see this fact from the decay ratio of the scalaron into a pair of light scalar particles during reheating epoch [8], [48]
-$\Gamma_{\phi\to\chi\chi}=\left(1-6\xi_{\chi}+\dfrac{2m_{\chi}^{2}}{M^{2}}\right)^{2}\,\dfrac{M^{3}\sqrt{1-4m_{\chi}^{2}/M^{2}}}{192\pi M_{p}^{2}}\,.$
-If $\xi_{\chi}=1/6$ and $m_{\chi}\to 0$, the decay ratio vanishes.
-In addition, if the mass of scalar field $\chi$ is $m_{\chi}>M/2$ where $M$ is scalaron mass term, the scalaron decay into a pair of $\chi$ particles is forbidden.
-Then, the perturbative production of $\chi$ particles is allowed only in the case of light scalars with mass $m_{\chi}<M/2$.
-However, it can also occur a scalar particle production by non perturbative effects by time-dependent background dynamics of $\phi$, $h$ and $a$ for a spectator field with mass $m_{\chi}>M/2$.
-In order to study this process it is necessary to find the equation of motion for the spectator field $\tilde{\chi}$ and track the mode evolution during and after inflation.
+In the paper [8], the authors study the gravitational production of dark matter in the $R^{2}$ inflation model and show that the scalaron $\phi$ can decay perturbatively in dark matter particles $\chi$ with masses $m_{\chi}<M/2$, where $M\simeq 1.3\times 10^{-5}M_{p}$ is the mass term of the scalaron.
+For heavy particles, perturbative particle production is kinematically forbidden.
+However, the production of scalar particles is not limited to perturbative methods; there may be particle production due to the expansion of the universe and the dynamics of the fields involved during the inflationary epoch.
+In our case, the classical time-dependent background is defined by the evolution of the scale factor $a(t)$ and the evolution of the scalar fields $\phi$ and $h$.
+That is, the matter-free scalar field $\chi$ propagates in a time-dependent classical background, whose dynamics are given by the evolution of the $\phi$ and $h$ fields of the Higgs-$R^{2}$ model.
+To study this process it is necessary to find the equation of motion for the spectator field $\tilde{\chi}$ and track the mode evolution during and after inflation.
 Varying the action (30) respect to $\tilde{\chi}$ yields the following equation of motion
-(□−$(\square-e^{-\alpha\phi}\,m_{\chi}^{2}-\xi_{\chi}\,R)\tilde{\chi}\\[5.69046pt] +\left[\dfrac{\alpha}{2}\,\left(1-6\xi_{\chi}\right)\square\phi-\dfrac{\alpha^{2}}{4}\left(1-6\xi_{\chi}\right)(\partial\phi)^{2}\right]\tilde{\chi}=0\,.$ (34)
+$(\square-e^{-\alpha\phi}\,m_{\chi}^{2}-\xi_{\chi}\,R)\tilde{\chi}+\left[\dfrac{\alpha}{2}\,\left(1-6\xi_{\chi}\right)\square\phi-\dfrac{\alpha^{2}}{4}\left(1-6\xi_{\chi}\right)(\partial\phi)^{2}\right]\tilde{\chi}=0\,.$ (32)
 Notice that the background dynamics is explicit in the equation of the spectator field $\tilde{\chi}$ due to the conformal transformation.
@@
 Consequently, in the Einstein frame, the model presents an interaction term with the dynamics of the classical background through the Ricci scalar, the scale factor $a$, and the dynamics of the fields $\phi$ and $h$.
-As a result, studying the solutions is generally highly complex from an analytical perspective; hence a numerical treatment of the background dynamics is necessary for a comprehensive analysis of mode evolution, specially after inflation when both fields $\phi$ and $h$ exhibit a strongly interacting anharmonic oscillatory behavior, as one can see in the figure 2.
+As a result, studying the solutions is generally highly complex from an analytical perspective; hence a numerical treatment of the background dynamics is necessary for a comprehensive analysis of mode evolution, especially after inflation when both fields $\phi$ and $h$ exhibit a strong interacting anharmonic oscillatory behavior, as one can see in the figure 2.
 By numerically determining the background evolution of the scalaron, the Higgs field, and the scale factor, we can compute the Ricci scalar using $R=6(\dot{H}+2H^{2})$, ensuring that the dynamics are well-determined.
 However, during effective inflation along one of the potential’s valleys, we can approximate the background dynamics using the single-field equation of motion (20).
-IV Mode evolution
-In this section, we numerically solve the equation of motion (34) along with the background equations (17), (18), (I) to obtain the mode functions $\chi$ for calculating the number density of particles produced, $n_{k}$.
-The equation of motion for $\tilde{\chi}$ considering a FLRW universe (I) can be written in terms of the cosmic time $t$ as
+4 Mode evolution
+In this section, we numerically solve the equation of motion (32) along with the background equations (16), (17), (1) to obtain the mode functions $\chi$ for calculating the number density of particles produced, $n_{k}$.
+The equation of motion for $\tilde{\chi}$ considering a FLRW universe (1) can be written in terms of the cosmic time $t$ as
 $\ddot{\tilde{\chi}}+3H\dot{\tilde{\chi}}+\Omega^{2}(t)\tilde{\chi}=0\,,$ (33)
@@
 We proceed by quantizing the spectator field in the Heisenberg picture so that the $\chi$ field and its conjugate momentum $\pi_{\chi}=\dot{\chi}$ are treated as field operators.
-In Fourier space the rescaled field can be written as
+In Fourier space, the rescaled field can be written as
 $X(t,\mathbf{r})=\int\dfrac{\mathrm{d}^{3}\mathrm{k}}{(2\pi)^{3}}\left(\hat{a}_{\mathbf{k}}\,X_{k}(t)\,e^{-i\mathbf{k}\cdot\mathbf{r}}+\hat{a}^{\dagger}_{\mathbf{k}}\,X^{*}_{k}(t)\,e^{i\mathbf{k}\cdot\mathbf{r}}\right)\,,$
@@
 and the following effective mass term
-$m_{\mathrm{eff}}^{2}(t)=e^{-\alpha\phi}m_{\chi}^{2}+\xi_{\chi}\,R-\dfrac{9}{4}H^{2}-\dfrac{3}{2}\dot{H}\\[5.69046pt] +3\alpha\left(\xi_{\chi}-\dfrac{1}{6}\right)(V_{\phi}-\alpha M_{p}^{2}\dot{H})\,.$ (42)
+$m_{\mathrm{eff}}^{2}(t)=e^{-\alpha\phi}m_{\chi}^{2}+\xi_{\chi}\,R-\dfrac{9}{4}H^{2}-\dfrac{3}{2}\dot{H}+3\alpha\left(\xi_{\chi}-\dfrac{1}{6}\right)(V_{\phi}-\alpha M_{p}^{2}\dot{H})\,.$ (40)
 The equation for $X_{k}$ describes an oscillator with time-dependent frequency $\omega_{k}^{2}(t)$, hence to track the evolution of the modes $X_{k}$ during and after inflation, it is necessary to analyze $\omega_{k}^{2}(t)$.
 One can see that during the quasi De Sitter phase during inflation when the effective single field inflation is valid, the effective mass is simplified because of $\dot{H}\to 0$.
-In this case one can see that the effective mass takes the form
-$m_{\mathrm{eff}}^{2}(t\ll t_{\mathrm{end}})\simeq e^{-\alpha\phi}m_{\chi}^{2}+12\left(\xi_{\chi}-\dfrac{3}{16}\right)H^{2}\\[5.69046pt] +3\alpha\left(\xi_{\chi}-\dfrac{1}{6}\right)V_{\phi}\,.$ (43)
+In this case, one can see that the effective mass takes the form
+$m_{\mathrm{eff}}^{2}(t\ll t_{\mathrm{end}})\simeq e^{-\alpha\phi}m_{\chi}^{2}+12\left(\xi_{\chi}-\dfrac{3}{16}\right)H^{2}+3\alpha\left(\xi_{\chi}-\dfrac{1}{6}\right)V_{\phi}\,.$ (41)
 We can see that the mass term is exponentially suppressed during inflation when $\phi\gg M_{p}$, making the mass of the spectator field much smaller than the Hubble parameter at the end of inflation, $m_{\chi}\ll H_{\mathrm{end}}$.
@@
 Additionally, if the coupling constant is $\xi_{\chi}<1/6$, the effective mass could be negative during inflation.
-In such a case, the frequency $\omega_{k}^{2}(t)$ can become negative for super-horizon modes $k\ll aH$, leading to a exponential mode growth $\tilde{\chi}\sim e^{|\omega_{k}|t}$ during inflation due to tachyonic instability.
+In such a case, the frequency $\omega_{k}^{2}(t)$ can become negative for super-horizon modes $k\ll aH$, leading to an exponential mode growth $\tilde{\chi}\sim e^{|\omega_{k}|t}$ during inflation due to tachyonic instability.
 This regime may result in a very efficient amplification of particle production [41, 33].
-Figure 3: Time-dependent frequency as a function of cosmic time t for a mass $m_{\chi}=10^{-2}M$, a non-minimal coupling $\xi_{\chi}=0$, and $k\gg aH$ for 5 e-folds before horizon crossing.
-We set $N_{e}=60$ before the end of inflation as the moment when the mode crosses the horizon.
 However, for large enough modes $k$, namely
@@
 For $t\gtrsim t_{\mathrm{end}}$, the frequency increases and oscillates after inflation as the background terms oscillate around the minimum of the potential (preheating).
+Figure 3: Time-dependent frequency as a function of cosmic time t for a mass $m_{\chi}=10^{-2}H_{\mathrm{end}}$, a non-minimal coupling $\xi_{\chi}=0$, and $k\gg aH$ for 5 e-folds before horizon crossing.
+We set $N_{e}=60$ before the end of inflation as the moment when the mode crosses the horizon.
 One can also observe growth due to the exponential dependence of the scalaron in the mass term toward the end of inflation.
-To numerically solve equation (40) along with the background equations, we must specify suitable intial conditions that ensure the modes are within the horizon at early times.
+To numerically solve equation (38) along with the background equations, we must specify suitable initial conditions that ensure the modes are within the horizon at early times.
 In the asymptotic limit $t\to 0$ and according to the behavior of $\omega_{k}^{2}(t)$ at the beginning of inflation, the dominant term is $k/a$.
@@
 […]
-V Description of particle production
-Once we have established the behavior of mode functions $X_{k}$ during and after inflation for some values of spectator mass, the non-minimal coupling $\xi_{\chi}$ and comoving wave number $k$, it is turn to analyze the non-perturbative particle production.
-The production of $\chi$ particles is due to the time dependence of the background variables, in particular the non-adiabatic behaviour of the frequency $\omega_{k}(t)$ that induces a mix among negative- and positive-frequency [3].
-The adiabaticity condition is given by the adiabatic paramater $A_{k}$ defined as
-Figure 4: An example of non-adiabatic mode evolution leading the gravitational particle production of $\chi$ particles described by the adiabaticity parameter $|A_{k}(t)|$ for $m_{\chi}=10^{-2}M$ and $\xi_{\chi}=1/6$.
-At the end of inflation, one see an abruptly increase on $|A_{k}(t)|$ and non-adiabaticity is induced, leading gravitational particle production $n_{k}(t)\neq 0$.
-At late times, when the stabilization occurs, we recover the adiabaticity evolution.
+5 Description of particle production
+Once we have established the behavior of mode functions $X_{k}$ during and after inflation for some values of spectator mass, the non-minimal coupling $\xi_{\chi}$ and comoving wavenumber $k$, it turns to analyze the non-perturbative particle production.
+The production of $\chi$ particles is due to the time dependence of the background variables, in particular, the non-adiabatic behavior of the frequency $\omega_{k}(t)$ that induces a mix among negative- and positive-frequency [3].
+The adiabaticity condition is given by the adiabatic parameter $A_{k}$ defined as
 $A_{k}(t)=\dfrac{\dot{\omega}_{k}}{\omega_{k}^{2}}\,.$ (44)
@@
 $\lim_{t\to 0}X_{k}(t)=\dfrac{e^{-i\int^{t}\omega_{k}(t^{\prime})\mathrm{d}t^{\prime}}}{\sqrt{2\omega_{k}(t)}}\,.$ (45)
-This behavior usually happens at early times $t_{i}\to 0$ and at sufficiently late times $t_{f}\to+\infty$, but for intermediate times $t_{i}<t_{*}<t_{f}$ during the expansion of the universe the evolution of the modes is no longer adiabatic $|A_{k}|\gg 1$.
+This behavior usually happens at early times $t_{i}\to 0$ and at sufficiently late times $t_{f}\to+\infty$, but for intermediate times $t_{i}<t_{*}<t_{f}$ during the early universe, the evolution of the modes is no longer adiabatic $|A_{k}|\gg 1$.
 During inflation, the positive-frequency solution corresponds to the Bunch-Davies vacuum (43), as described in the previous section.
@@
 From figure 3, we can see that the evolution of $\omega_{k}(t)$ is smooth until the end of inflation.
-At that point, the condition $|A_{k}|\ll 1$ is strongly violated, leading to an enhancement in particle production and gravitational particle production is enhanced at the time $t_{*}$, which corresponds to the end of inflation.
+At that point, the condition $|A_{k}|\ll 1$ is strongly violated, leading to an enhancement in particle production, and gravitational particle production is enhanced at the time $t_{*}$, corresponding to the end of inflation.
 We show the non-adiabaticity evolution in the figure 4 parametrized by the adiabatic parameter $|A_{k}(t)|$.
@@
 Comparing with the Bunch-Davies initial condition (45), one see thats $\alpha_{k}(t_{0})=1$ and $\beta_{k}(t_{0})=0$.
-The comoving particle number density $n_{\chi}(t)$ is related to the coefficient $\beta_{k}$ of the Bogolyubov transformations which relate the vacuum states at the asymptotic limits $t\to 0$ and $t\to\infty$, and coincides with $\beta_{k}(t)$ at late times.
+Figure 4: An example of non-adiabatic mode evolution leading the gravitational particle production of $\chi$ particles described by the adiabaticity parameter $|A_{k}(t)|$ for $m_{\chi}=10^{-2}H_{\mathrm{end}}$ and $\xi_{\chi}=1/6$.
+At the end of inflation, one sees an abrupt increase on $|A_{k}(t)|$ and non-adiabaticity is induced, leading to gravitational particle production $n_{k}(t)\neq 0$.
+At late times, when the stabilization occurs, we recover the adiabaticity evolution.
+In other words, gravitational particle production occurs at very early times $t_{\mathrm{in}}\to 0$ and at late times $t_{f}\to\infty$ if the initial and final vacuum states are not the same.
+We define the initial vacuum state with the corresponding creation-annihilation operators $\hat{a}^{\dagger}_{\mathbf{k}}$ and $\hat{a}_{\mathbf{k}}$ by
+$\hat{a}_{\mathbf{k}}\left|0\right>_{\mathrm{early}}=0\,.$ (47)
+The final state will have a different set of creationist-annihilation operators, and thus a different vacuum state.
+$\hat{a}_{\mathbf{k}}$ it is related to the late-time operators $\hat{b}^{\dagger}_{\mathbf{k}}$ and $\hat{b}_{\mathbf{k}}$ by a Bogolyubov tranformation
+$\hat{a}_{\mathbf{k}}=\alpha^{*}_{\mathrm{k}}\hat{b}_{\mathbf{k}}-\beta^{*}_{\mathrm{k}}\hat{b}^{\dagger}_{\mathrm{k}}\,,$ (48)
+where $\alpha_{\mathbf{k}}$ and $\beta_{\mathbf{k}}$ are the time-independent Bogolyubov coefficients.
+The comoving particle number density $n_{\chi}(t)$ is related to the coefficient $\beta_{k}$ of the Bogolyubov transformations which relate the vacuum states at the asymptotic limits $t\to 0$ and $t\to\infty$ and coincides with $\beta_{k}(t)$ at late times.
 We can compute $\beta_{k}(t)$ by solving (38) and evaluating
 $|\beta_{k}(t)|^{2}=\dfrac{\omega_{k}}{2}|X_{k}|^{2}+\dfrac{|\dot{X}_{k}|^{2}}{2\omega_{k}}-\dfrac{1}{2}\,.$ (49)
-Figure 5: Left: Comoving number density of particles created for $m_{\chi}=10^{-1}M$ and $\xi_{\chi}=1/6$.
-The graph shows the evolution of the occupation number $n_{k}/k_{\mathrm{end}}^{3}$ as a function of time $t$, where the mode $k$ was chosen to be 5 e-folds before leaving the horizon during inflation and $k_{\mathrm{end}}=a_{\mathrm{end}}H_{\mathrm{end}}$ is the mode at the end of inflation.
-Right: Spectrum of the gravitationally produced scalar particles for the same numerical example.
-Modes with $k=aH$ are on the horizon scale at the end of inflation.
-The mass and the momenta $k$ both are given in units of $M$.
 The physical number density $n_{\chi}(t)$ in the late-time limit $t\to\infty$ is given by
@@
 where $n_{k}(t)$ is the comoving number density of particles with comoving wavenumber $k$.
-V.1 Numerical results
-What follows is to find the spectrum of $\chi$ particles and its evolution in the cosmic time $t$.
-Hence we numerically solve (40), constructed the Bogolyubov coeffcient $|\beta_{k}(t)|$ and compute the spectrum $n_{k}$ using (50).
-In the previous section we described the procedure to numerically evaluate the set of equations to ensure that each mode $k$ is within the horizon at the beginning of the evolution, which allows us to use the Bunch-Davies vacuum (45) as a initial condition of the mode equation for $X_{K}(t)$.
+Notice that the expression (49) is only valid when $\omega_{k}^{2}(t)>0$, so the comoving number density of particles created is valid at times when the frequency (39) is positive.
+5.1 Numerical results
+What follows is to find the spectrum of $\chi$ particles and their evolution in the cosmic time $t$.
+Hence we numerically solve (38), constructed the Bogolyubov coefficient $|\beta_{k}(t)|$, and compute the spectrum $n_{k}$ using (50).
+In the previous section, we described the procedure to numerically evaluate the set of equations to ensure that each mode $k$ is within the horizon at the beginning of the evolution, which allows us to use the Bunch-Davies vacuum (43) as an initial condition of the mode equation for $X_{K}(t)$.
 The initial conditions for the background are set at time $t_{\mathrm{ini}}$, corresponding to the moment when the modes are within the horizon.
 For example, for $N_{*}=60$ at the time of crossing, the initial time corresponds to $t_{\mathrm{ini}}\approx 28.712$ in units of $[10^{-5}M_{p}^{-1}]$, and the corresponding mode $k=aH$ will be the value of $aH$ at time $t_{\mathrm{ini}}$.
-Then, the Bogolyubov coefficient $|\beta_{k}(t)|^{2}$ is calculated using (49) and the comoving number density of gravitational particle production is determined by $n_{k}(t)$.
-V.1.1 Conformal coupling $\xi_{\chi}=1/6$
-First of all, we focus on the conformal coupling $\xi_{\chi}=1/6$ and consider light masses $m_{\chi}\ll M/2$.
-As first example, we show the the comoving number density of particles $n_{k}(t)$ as a function of time $t$ and the corresponding spectrum for the case of $m_{\chi}=10^{-1}M$.
-In the left panel of figure 5, we observe that the number of particles created saturates at the end of inflation, stabilizing and becoming constant at late times $t\to\infty$.
-This growth in $n_{k}$ is due to the non-adiabaticity of $\omega_{k}$ at the end of inflation, as can be seen in the figure 4.
-The oscillations of $n_{k}(t)$ after inflation are due to resonances at certain moments between the oscillations of the modes $X_{k}$ and the frequency itself.
-In the right panel, we observe the corresponding spectrum of $n_{k}(t)$.
-Several remarkable features can be observed.
-For IR modes $(k\ll aH)$, we see that the spectrum is blue-tilted, reaching a peak at some intermediate value $k_{\mathrm{peak}}$, and then is suppressed for UV modes $(k\gg aH)$ due to inflation, that is, those that remain within the horizon during inflation.
-Modes with $k\ll k_{\mathrm{end}}$ leaves the horizon during inflation and re-enter later during the reheating epoch after inflation.
+Then, the Bogolyubov coefficient $|\beta_{k}(t)|^{2}$ is calculated using (49), and the comoving number density of gravitational particle production is determined by $n_{k}(t)$.
+5.1.1 Conformal coupling $\xi_{\chi}=1/6$
+We first present the results for conformally coupled particles ($\xi_{\chi}=1/6$).
+As an example, we present the time evolution of the comoving number of particles created $n_{k}(t)$ as a function of time for two different cases: light particles $m_{\chi}<H_{\mathrm{end}}$ and heavy particles $m_{\chi}\gtrsim H_{\mathrm{end}}$.
+We should note that the expression for $\beta_{k}(t)$ is only valid when $\omega_{k}^{2}(t)>0$ , we point out that $n_{k}(t)$ only makes physical sense at late times $t\to\infty$ [5].
+The frequency (39) for scalar conformally coupled turns out to be
+$\omega_{k}^{2}(t)=\left(\dfrac{k}{a}\right)^{2}+e^{-\alpha\phi}m_{\chi}^{2}+\dfrac{R}{6}-\dfrac{9}{4}H^{2}-\dfrac{3}{2}\dot{H}$ (51)
+which can be negative for $k\ll aH$.
+For this reason, we take into account the sub-horizon evolution of the modes $X_{k}(t)$ and we take $k$ deep inside the horizon at the beginning of the simulation.
+In the figure, we present two cases: $m_{\chi}=10^{-5}H_{\mathrm{end}}$ and $m_{\chi}=10^{-1}H_{\mathrm{end}}$.
+In both panels of figure 5, we observe that the number of particles created saturates at the end of inflation, stabilizing and becoming constant at late times $t\to\infty$.
+This growth in $n_{k}$ is due to the non-adiabaticity of $\omega_{k}$ at the end of inflation, as can be seen in figure 4.
+Figure 5: Comoving number density of particles, plotted on a logarithmic scale, as a function of cosmic time $t$ for two different values of the spectator mass $m_{\chi}$, with the same mode $k$, corresponding to $N_{k}=60$.
+The graph depicts the evolution of the occupation number $n_{k}/k_{\mathrm{end}}^{3}$, where $k_{\mathrm{end}}=a_{\mathrm{end}}H_{\mathrm{end}}$ denotes the mode at the end of inflation
+The oscillations of $n_{k}(t)$ after inflation arise from resonances occurring at specific moments between the oscillations of the modes $X_{k}$ and the frequency itself.
+During this epoch, particle production is driven primarily by the oscillatory behavior of the scalaron and the Higgs field, as illustrated in figure 2, through the term $\dot{H}$ and the Ricci scalar $R$ via the non-minimal coupling $\xi_{\chi}=1/6$.
+Additionally, we observe that particle production near the end of inflation is most efficient for light particles.
+In both cases, the particle number is well-defined, grows from the end of inflation, and stabilizes at late times, as expected.
+Figure 6: Numerical $n_{k}/k_{\mathrm{end}}^{3}$ vs $k/k_{\mathrm{end}}$ spectra of conformally-coupled scalars for differente light masses $m_{\chi}$.
+One can see that the UV tail is suppressed for this range of masses.
+Figure 7: Right: Comoving number density of particles as a function fo cosmic time $t$.
+Left: Numerical $n_{k}/k_{\mathrm{end}}^{3}$ vs $k/k_{\mathrm{end}}$ spectra of conformally-coupled scalars for differente large masses $m_{\chi}\gtrsim H_{\mathrm{end}}$.
+One can see that the UV tail is suppressed for this range of masses.
+We have numerically computed the spectra of the comoving number density $n_{k}(t)$ as a function of $k/k_{\mathrm{end}}$ for a range of light masses varying from $10^{-5}H_{\mathrm{end}}$ to $10^{-1}H_{\mathrm{end}}$.
+Our findings are summarized in figure 6.
+From the figure, one can observe several remarkable features.
+Conformally coupled scalars show a spectra that are strongly peaked at some value $k_{\mathrm{peak}}$.
+For IR modes $(k\ll k_{\mathrm{end}})$, we see that the spectrum is blue-tilted, reaching a peak at some intermediate value $k_{\mathrm{peak}}>k_{\mathrm{end}}$, and then is suppressed for UV modes $(k\gg k_{\mathrm{end}})$ due to inflation, that is, those that remain within the horizon during inflation.
+IR modes leave the horizon during inflation and re-enter later during the reheating epoch after inflation.
 Therefore, the dynamics of the $X_{k}(t)$ modes are dictated by the effective mass term (41) with a quasi-De Sitter spacetime as background with $H\simeq\mathrm{constant}$.
-It is important to note that the UV modes cross the horizon just at the end of inflation, leading to the production of these modes at that instant and their subsequent evolution during the preheating epoch after inflation, as seen in the left panel of figure 5.
-During this epoch, the production occurs mainly due to the oscillatory nature of the scalaron and the Higgs field, as shown in figure 2, through the term $\dot{H}$ and the Ricci scalar $R$ via the non-minimal coupling $\xi_{\chi}=1/6$.
+It is important to note that the UV modes do not cross the horizon during inflation.
+In place, these modes are mainly produced during the transition to the matter-dominated era and in general during the reheating era.
 The suppression of UV modes occurs because the $\chi$ scalars become supermassive at the end of inflation due to the exponentially scalaron-dependence mass, causing the effective mass to become substantial.
 As a result, gravitational particle production becomes less efficient, leading to a significant suppression in the particle density number $n_{k}(t)$.
-Figure 6 presents the spectrum of the comoving number of particles produced across the range of light masses considered in this study.
-Figure 6: Comoving spectrum of gravitational scalar particle production $n_{k}$ for a range of light masses of the field $\chi$ as a function of rescaled horizon modes $k/k_{\mathrm{end}}$.
-Masses are given in units of scalaron mass $M$.
-One can seen that the UV tail is suppressed for light masses $m_{\chi}<M/2$.
-The numerical results exhibit a blue-tilted behavior in the infrared (IR) modes, with an exponential suppression in the ultraviolet (UV) tail for light masses $m_{\chi}<M/2$.
 The mass of the field plays a key role in determining the amplitude of $n_{k}$ at the peak momentum $k_{\mathrm{peak}}$, which systematically decreases as $m_{\chi}$ increases [43].
-It is important to emphasize that, in this mass range, particle production following inflation is predominantly driven by the decay of the scalaron into pairs of light $\chi$ particles, as described by expression (33).
-Furthermore, figure 6 indicates that non-perturbative particle production within this mass range remains minimal, suggesting that such production could lead to an overproduction [8, 50].
-For the specific cases of $m_{\chi}=0.2M$ and $m_{\chi}=0.3M$, the production is significantly more moderate.
-The spectra also exhibit a tendency to flatten in the UV region, implying that gravitational particle production becomes increasingly inefficient for larger masses.
-V.1.2 Non-minimal coupling $\xi_{\chi}=0$
-Figure 7: Comoving spectrum of gravitational scalar particle production $n_{k}$ for $\xi_{\chi}=0$ as a function of rescaled horizon modes $k/k_{\mathrm{end}}$.
-Masses are given in units of scalaron mass $M$.
-One can seen that the spectrum is nearly scale-invariant with a blue-tilted behavior for larger masses.
-Figure 8: Comoving number density of particles $n_{k}/k_{\mathrm{end}}^{3}$ for $m_{\chi}=10^{-1}M$ and $\xi_{\chi}=0$ as a function of cosmic time $t$.
-For the case of $\xi_{\chi}=0$, we have a different situation. During inflation, the effective mass (43) can be negative for light masses $m_{\chi}\ll H_{\mathrm{end}}$
-$m_{\mathrm{eff}}^{2}(t)\approx e^{-\alpha\phi}m_{\chi}^{2}-\dfrac{9}{4}H^{2}-\dfrac{\alpha}{2}V_{\phi}<0\,,$
-and once the modes cross the horizon, we obtain $\omega_{k}^{2}(t)<0$, leading to tachyonic instabilities as pointed out in the discussion following equation (43).
-Therefore, particle production can be due to resonances from oscillations or due to the instabilities caused by $\omega_{k}^{2}(t)<0$.
-In figure 8, we can observe the behavior of $n_{k}(t)$ normalized to its value at the end of inflation as a function of cosmic time.
-We can see that for the same mass value $m_{\chi}=10^{-1}M$, particle production is significantly larger than in the case of conformal coupling (see figure 5).
-The tachyonic growth of the modes $X_{k}(t)$ outside the horizon is the main driver for the enhanced production of minimally coupled particles compared to the $\xi_{\chi}=1/6$ case.
-We observe that after inflation, $n_{k}(t)$ exhibits an anharmonic oscillatory behavior with maximum production peaks until it reaches a constant value as $t\to\infty$.
-However, for masses $m_{\chi}\gg H_{\mathrm{end}}$, the dominant terms in the effective mass during inflation are $e^{-\alpha\phi}m_{\chi}^{2}$ and $(k/a)^{2}$, which implies a positive frequency $\omega_{k}^{2}(t)>0$.
-In figure 7, we show the spectrum of scalar particles created in the case of minimal coupling and masses $m_{\chi}\gtrsim M/2$.
-As can be seen, the spectrum generated in the IR region is qualitatively different as the mass increases.
-We observe an almost scale-invariant spectrum that is slightly blue-tilted for masses $m_{\chi}>M/2$.
-VI Discussion and remarks
+For large masses $m_{\chi}\gtrsim H_{\mathrm{end}}$, the situation is quite similar.
+As the mass increases, the particle production tends to decrease due to the increase of the effective mass.
+This can be seen in the left panel of figure 7 for a mass of $m_{\chi}=H_{\mathrm{end}}$ and for a mode crossing the horizon at $N_{k}=60$.
+In the rigt panel, we present the comoving spectra of particle production for $m_{\chi}\gtrsim H_{\mathrm{end}}$.
+We find that the behaviour for UV modes is different than in the case of light masses.
+The location of the peak $k_{\mathrm{peak}}$ is different as a function of the mass value, once $k_{\mathrm{end}}$.
+While $k_{\mathrm{peak}}>k_{\mathrm{end}}$ as in the case of light particles, the entry to the horizon, and hence the tail suppression of the UV modes is different.
+This is largely due to the possible resonances that exist for large masses, as the mass term $e^{-\alpha\phi}$ dominates after inflation when $\phi<M_{p}$, which can excite different modes after inflation.
+Figure 8: The comoving spectrum of gravitational scalar particle production $n_{k}$ for $\xi_{\chi}=0$ as a function of rescaled horizon modes $k/k_{\mathrm{end}}$.
+One can see that the spectrum is nearly scale-invariant for masses $m_{\chi}\lesssim H_{\mathrm{end}}$ and blue-tilted behavior for larger masses
+5.1.2 Minimal coupling $\xi_{\chi}=0$
+For the case of $\xi_{\chi}=0$, we have a different situation. During inflation, the frequency (39) is given by
+$\omega_{k}^{2}(t)=\left(\dfrac{k}{a}\right)^{2}+e^{-\alpha\phi}m_{\chi}^{2}-\dfrac{9}{4}H^{2}+\dfrac{7}{6}\dot{H}-\dfrac{\alpha}{2}V_{\phi}\,.$ (52)
+In this case, the derivative potential $V_{\phi}$ and the Hubble parameter $H^{2}$ contribute so that the effective mass $m_{\mathrm{end}}$ is negative.
+Thus, once the modes cross the horizon and for masses $m_{\chi}\ll H_{\mathrm{end}}$, this may lead to $\omega_{k}^{2}(t)<0$, resulting in tachyonic instabilities, as mentioned in the discussion following equation (41) [41].
+These instabilities primarily affect the IR modes of the produced particle spectrum, as the modes outside the horizon during inflation experience tachyonic growth.
+This growth of the modes $X_{\mathbf{k}}(t)$ outside the horizon is the main driver of the enhanced particle production for minimally coupled scalars compared to conformally coupled scalars.
+The spectrum of the produced particles is also affected, and its behavior depends on the mass $m_{\chi}$.
+To obtain the gravitational spectra, we use the same technique to solve the equation numerically: we choose the $k$-mode inside the horizon, approximately 5 e-folds before it crosses the horizon and show the numerical result for $\beta_{k}(t)$ at those times when $\omega_{k}^{2}>0$.
+In the figure 8, we show the spectrum of created particles for minimally coupled masses.
+We observe that the spectrum is nearly scale-invariant for IR modes and for masses $m_{\chi}\lesssim H_{\mathrm{end}}$ before reaching the peak at $k_{\mathrm{peak}}$.
+However, as the mass increases, the spectrum tends to be blue-tilted.
+In the UV region, all modes decay either as a power law or exponentially.
+Figure 9: Comoving number density of particles as a function of cosmic time.
+A similar behavior can be observed in Figure 6: the number of created particles saturates at the end of inflation and stabilizes, becoming constant at late times
+A clear difference exists between the cases $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$.
+In the frequency $\omega_{k}^{2}(t)$, the Ricci scalar does not appear for $\xi_{\chi}=0$, indicating a significant difference in the behavior of $\omega_{k}$ for different coupling cases.
+This is because the Ricci scalar exhibits an oscillatory behavior after inflation due to the oscillations of the fields $\phi$ and $h$.
+In other words, for the case $\xi_{\chi}=1/6$, the Ricci scalar plays an important role in describing the evolution of the modes $X_{k}(t)$, which translates into a different spectrum of created particles.
+From figure 9 we can see that the particle production at minimal coupling is much larger than the conformally-coupled case.
+This is mainly due to the behavior of the spectator field $X_{k}(t)$ during inflation.
+As explained above, the tachyonic instabilities in the case of $\xi_{\chi}=0$ cause the modes $X_{k}(t)$ to grow exponentially.
+This growth of the super-horizon modes is primarily responsible for the enhanced gravitational particle production for minimally coupled scalars.
+6 Discussion and remarks
 In this study, we investigated the gravitational and non-perturbative production of free scalar particles, $\chi$, in Higgs-$R^{2}$ inflation.
-We began with a free spectator field, $\chi$, interacting solely through the gravitational term $\xi_{\chi}R$ in the Jordan frame.
-By applying a conformal transformation to transition into the Einstein frame, direct couplings between the $\chi$ field and the scalaron $\phi$ emerged.
-Within the context of inflationary dynamics in the Einstein frame, we derived the equation of motion for the $\chi$ field, taking into account the non-minimal coupling with R.
-We numerically solved this system of equations, imposing Bunch-Davies conditions on the $\chi$ field 5 e-folds before the modes crossed the horizon, to determine the mode functions of the $\chi$ field.
-From these solutions, we extracted the Bogolyubov coefficient $\beta_{k}(t)$ to calculate the comoving particle density and visualize their spectrum.
-We explored two distinct scenarios for the non-minimal coupling $\xi_{\chi}$, considering different mass ranges.
-For $\xi_{\chi}=1/6$ and light masses, we observed efficient particle production at the end of inflation due to the term $e^{-\alpha\phi}m_{\chi}^{2}$, which becomes dominant at that stage.
-The resulting power spectrum exhibited a blue-tilted behavior, indicating that particle production was dominated by modes near the Hubble radius $(aH)^{-1}$ at the end of inflation, while shorter wavelength modes were suppressed.
-In the case of $\xi_{\chi}=1/6$ and for masses close to $m_{\chi}\sim M/2$, the spectrum tends to flatten out without exhibiting a peak in the UV region.
-This phenomenon is possibly due to the mass term that depends exponentially on the scalaron $\phi$, causing the effective mass to become superheavy at the end of inflation.
-In contrast, for $\xi_{\chi}=0$, the spectrum remained nearly scale-invariant for masses $m_{\chi}\gtrsim M/2$, with a blue tilt emerging for larger masses.
-For light masses and $\xi_{\chi}=0$, we find that the comoving number of particles created is significantly higher compared to the case of conformally coupled particles, which is due to the tachyonic growth of the modes $X_{k}(t)$.
-We have not attempted a detailed exploration of the parameter space of the Higgs-$R^{2}$ inflation model or the parameters of the scalar field $\chi$.
-Instead, we have taken the Higgs-$R^{2}$ inflation model, whose observables are consistent with Planck, and considered the possibility of non-perturbative particle production through the direct couplings that arise between the $\chi$ field and the inflationary sector when changing the conformal frame.
-We found that inflation driven by the Higgs field and the scalaron can produce scalar particles in the cases $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$ and for a specific range of masses.
-This field could potentially contribute to the abundance of dark matter, for which its current abundance would need to be calculated.
-There remains much to be done for future work, such as considering the production of spin-$1/2$ and spin-$1$ dark matter, the production of gauge bosons $W^{\pm}_{\mu}$, $Z_{\mu}$, as well as the implications on primordial non-Gaussianities in the CMB induced by particle production.
+We started with a free spectator field, $\chi$, interacting only through the gravitational term $\xi_{\chi}R$ in the Jordan frame.
+By applying a conformal transformation to transition into the Einstein frame, direct couplings between the $\chi$ field and the scalaron $\phi$ naturally emerged.
+Within the context of inflationary dynamics in the Einstein frame, we derived the equation of motion for the $\chi$ field, accounting for its non-minimal coupling to the scalar curvature $R$.
+We numerically solved this system of equations, imposing Bunch-Davies conditions on the $\chi$ field five e-folds before the modes crossed the horizon, to determine the mode functions of $\chi$.
+From these solutions, we extracted the Bogolyubov coefficient $\beta_{k}(t)$ to compute the comoving particle density and visualize the resulting spectrum.
+We explored two distinct scenarios for the non-minimal coupling parameter $\xi_{\chi}$, considering different mass ranges.
+For $\xi_{\chi}=1/6$, we observed efficient particle production at the end of inflation, driven by the term $e^{-\alpha\phi}m_{\chi}^{2}$, which becomes dominant at this stage.
+The resulting power spectrum exhibited a blue-tilted behavior, indicating that particle production was dominated by modes near the Hubble radius $(aH)^{-1}$ at the end of inflation, while shorter-wavelength modes were suppressed.
+As the mass increased in the $\xi_{\chi}=1/6$ case, particle production tended to decrease due to the increase in the effective mass.
+In contrast, for $\xi_{\chi}=0$, the spectrum remained nearly scale-invariant for masses $m_{\chi}\lesssim H_{\mathrm{end}}$, with a blue tilt emerging for larger masses $m_{\chi}\gtrsim H_{\mathrm{end}}$.
+For light masses and $\xi_{\chi}=0$, we found that the comoving number of particles created was significantly higher compared to the case of conformally coupled particles, primarily due to the tachyonic growth of the modes $X_{k}(t)$.
+We did not attempt a detailed exploration of the parameter space of the Higgs-$R^{2}$ inflation model or the parameters of the scalar field $\chi$.
+Instead, we focused on the Higgs-$R^{2}$ inflation model, whose observables are consistent with Planck data, and considered the possibility of non-perturbative particle production through the direct couplings that arise between the $\chi$ field and the inflationary sector when changing the conformal frame.
+We found that inflation driven by the Higgs field and the scalaron can produce scalar particles in the cases $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$ for a specific range of masses.
+This field could potentially contribute to the abundance of dark matter, though its present-day abundance would need to be calculated.
+There remain several directions for future research, including the production of spin-$1/2$ and spin-$1$ dark matter, the production of gauge bosons $W^{\pm}_{\mu}$, $Z_{\mu}$, and the implications of particle production on primordial non-Gaussianities in the CMB.
 It would also be interesting to study the isocurvature constraints for light scalar field masses in this model.
-We leave these intriguing topics for future studies.
-VII Acknowledgements
+We leave these intriguing topics for future investigation.
+[…]
+Acknowledgements
 We are very grateful to Dr.
@@
 has been sponsored by CONACHyT-Mexico through a doctoral scholarship.
-This work was also partially supported by CONACHyT-Mexico through the support CBF2023-2024-1937.
+This work was also partially supported by CONACHyT- Mexico through the support CBF2023-2024-193

### v3

@@
 For both scenarios, we consider light particles with masses $m_{\chi}\ll H_{\mathrm{end}}$ and large masses that exceed the Hubble scale at the end of inflation $m_{\chi}\gtrsim H_{\mathrm{end}}$.
+We use these numerical results to calculate the relic abundance $\Omega_{\chi}h^{2}$ to find out if the model is viable as a dark matter candidate.
 keywords:
@@
 This oscillatory behavior enhances scalar particle production through non-perturbative mechanisms such as tachyonic instabilities [32, 33].
+The preheating phase after inflation has been studied [34, 35, 36, 37], as well as the production of primordial magnetic fields during the inflationary phase [38], considering effective single-field inflation.
 We propose an extension to the Higgs-$R^{2}$ model by introducing a massive scalar field $\chi$ as a spectator field that does not contribute to the energy density during the inflationary epoch, with negligible interactions with the SM but considering a direct coupling to the Ricci scalar via a non-minimal coupling term $\xi_{\chi}R\,\chi^{2}$.
@@
 The article is organized as follows: Section 2 provides a brief introduction to the Higgs-$R^{2}$ model and the background dynamics during and after inflation for a set of suitable initial conditions.
-Section 3 analyzes the couplings induced by gravitational effects in the Einstein frame, in Section 4, we derive the mode equation for the spectator field, which is necessary to investigate the behavior of the time-dependent frequency during and after inflation.
+Section 3 analyzes the couplings induced by gravitational effects in the Einstein frame.
+Section 4, we derive the mode equation for the spectator field, which is necessary to investigate the behavior of the time-dependent frequency during and after inflation.
 We adopt a numerical approach to solve the mode equation for modes that leave 5 e-folds before the horizon crossing during inflation, ensuring that each mode is within the horizon at the start of the numerical simulation.
 Section 4 examines the production of massive particles during preheating.
-Finally, Section 5 discusses the results and conclusions.
+In section 5 we show the numerical results for the gravitational production spectra for $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$ and for the time-evolution of $\beta_{k}$-Bogolyubov coefficient.
+Lastly, section 6 is dedicated to the relic abundance of gravitational production of scalars.
 In the following, we will use natural units in which $\hbar=1$, $c=1$, $\kappa_{\mathrm{B}}=1$ and we use the mass Planck as $8\pi G=M_{p}^{-2}$.
@@
 To describe the inflationary dynamics of the model, it is more convenient to bring the action (2) into the canonical Einstein-Hilbert form.
-So we need to carry out a conformal transformation $\Omega^{2}(x)$ introducing an extra scalar degree of freedom $\phi$ called scalaron, which comes from the quadratic term in $R_{J}$.
+So we need to carry out a conformal transformation $\Psi^{2}(x)$ introducing an extra scalar degree of freedom $\phi$ called scalaron, which comes from the quadratic term in $R_{J}$.
 Thus, to go from the Jordan frame to Einstein frame one has to perform a conformal transformation of the space-time metric $g_{\mu\nu}\to\Psi^{2}(x)g_{\mu\nu}$ with $\Psi(x)>0$, where the correct form of it is
@@
 Then in the Einstein frame, the Higgs-$R^{2}$ model of inflation takes the form of a two-scalar fields model with a non-canonical kinetic term
-$S_{E}=\int\mathrm{d}^{4}x\,\sqrt{-g}\,\left\{\dfrac{M_{p}^{2}}{2}R-\dfrac{1}{2}g^{\mu\nu}\partial_{\mu}\phi\,\partial_{\nu}\phi-\dfrac{1}{2}\,e^{-\alpha\phi}g^{\mu\nu}\partial_{h}\,\partial_{\nu}h-V(\phi\,,h)\right\}\,,$ (5)
-which can be written in the general action for multi-field inflationary models
 Figure 1: Scalar potential in the Einstein frame of the Higgs-$R^{2}$ model seen from different angles.
@@
 We have restricted ourselves to the standard parameter space $\xi_{h}=4000$, $\xi_{s}=2\times 10^{9}$ and $\lambda=0.013$.
+$S_{E}=\int\mathrm{d}^{4}x\,\sqrt{-g}\,\left\{\dfrac{M_{p}^{2}}{2}R-\dfrac{1}{2}g^{\mu\nu}\partial_{\mu}\phi\,\partial_{\nu}\phi-\dfrac{1}{2}\,e^{-\alpha\phi}g^{\mu\nu}\partial_{h}\,\partial_{\nu}h-V(\phi\,,h)\right\}\,,$ (5)
+which can be written in the general action for multi-field inflationary models
 $S=\int\mathrm{d}^{4}x\,\sqrt{-g}\left[\dfrac{M_{p}^{2}}{2}R-\dfrac{1}{2}G_{IJ}\,g^{\mu\nu}\partial_{\mu}\phi^{I}\,\partial_{\nu}\phi^{J}-V(\phi^{I})\right]\,,$ (6)
@@
     $V(\phi\,,h\ll M_{p})\simeq\dfrac{M_{p}^{4}}{4\xi_{s}}(1-e^{-\alpha\phi})^{2}\,.$ (11)
-We can see the positions of the valleys and the hill at $h=0$ in figure 1.
+We can see the positions of the valleys and the hill at $h=0$ in Fig.
+1.
 The inflationary dynamics occur along one of these valleys, showing a universal attractor behavior, so that inevitably the fields $\phi$ and $h$ will always fall into one of the valleys and then drive slow-roll inflation.
@@
 The inflation occurs when $\phi\gg M_{p}$ and the value of the auxiliary field $h$ is given by (13) inside one of the valleys of the potential.
-In the figure 2 we present the oscillatory regime of both fields after inflation in terms of cosmic time for the initial conditions $\phi_{0}=5.7M_{p}$ and $h_{0}\approx 0.096M_{p}$.
+In the Fig.
+2 we present the oscillatory regime of both fields after inflation in terms of cosmic time for the initial conditions $\phi_{0}=5.7M_{p}$ and $h_{0}\approx 0.096M_{p}$.
 However, according to the argument we have presented above, during inflation the isocurvature effects are absent and the dynamics inside the valleys are an effective single-field description.
-Then the scalaron field should fulfill the single-field equation
+Then the scalaron field should fulfill the single-field equation [14, 38]
 $\ddot{\phi}+3H\dot{\phi}+\dfrac{\mathrm{d}V_{\mathrm{eff}}}{\mathrm{d}\phi}\simeq 0\,,$ (20)
 and the evolution of the Higgs field is determined by the condition (13).
-We can assume a single-field slow roll approximation in the inflationary dynamics inside the valleys.
-This happens when $\ddot{\phi}\ll 3H\dot{\phi}$, thus the field equation for $\phi$ is
-$3H\dot{\phi}+\dfrac{\mathrm{d}V_{\mathrm{eff}}}{\mathrm{d}\phi}\simeq 0\,,$ (21)
-which can be integrated by quadrature together with the Friedman equations
-$\phi(t)=\phi_{0}+\alpha^{-1}\log\left(1-\dfrac{t}{t_{\mathrm{end}}}\right)\,,$ (22)
-$a(t)=e^{H_{0}t}\,\left(1-\dfrac{t}{t_{\mathrm{end}}}\right)^{3/4}\,,$ (23)
-$H(t)=H_{0}\,-\dfrac{3}{2}\,t_{\mathrm{end}}^{-1}\,\left(1-\dfrac{t}{t_{\mathrm{end}}}\right)^{-1}\,,$ (24)
-where
-$t_{\mathrm{end}}=\dfrac{3}{4}\,H_{0}^{-1}\,e^{\alpha\phi_{0}}\,,\quad H_{0}\simeq\dfrac{M_{p}^{4}}{\sqrt{12}(\xi_{s}+\xi_{h}^{2}/\lambda)}\,,$ (25)
-is the time that the effective inflation ends and $H_{0}$ is the scale Hubble parameter during inflation, with $\phi_{0}\sim 5.7M_{p}$ is the value of $\phi$ at the beginning of inflation.
-It is clear from slow-roll solutions that during inflation we have the usual quasi-De Sitter solutions for $\phi$, $a$, and $H$.
-The effective single-field inflation ends when the slow roll parameter $\epsilon$ is $\epsilon(t_{\mathrm{end}})=1$, which happens at the final values
-$\phi(t_{\mathrm{end}})$ $=$ $\alpha^{-1}\log\left(1+\dfrac{2}{\sqrt{3}}\right)\approx 0.92M_{p}\,,$ (26)
-$H_{e}$$\approx$ $0.54\,H_{0}\,.$ (27)
+It is worth mentioning that the effective single-field approximation is only valid during inflation within one of the valleys of the potential (8).
+The effective single-field inflation ends when the slow roll parameter $\epsilon$ is $\epsilon(t_{\mathrm{end}})=1$, which happens at the final value $\phi(t_{\mathrm{end)}}\approx 0.92M_{p}$, where
+$t_{\mathrm{end}}=\dfrac{3}{4}\,H_{0}^{-1}\,e^{\alpha\phi_{0}}\,.$ (21)
+After inflation, during the reheating era, the scalaron $\phi$ and the Higgs field $h$ oscillate around the global minimum $(0,0)$ of (8).
+That is, to follow the complete evolution of the fields after inflation, the system of equations (16), (17), (18), and (19) must be solved numerically with the appropriate initial conditions.
+Inflation within the valleys occurs for super-Planckian values of the field $\phi\gg M_{p}$ in order to produce sufficient inflation, i.e., $N_{\mathrm{end}}\gtrsim 60$.
+This condition is satisfied if $\phi_{0}\simeq 5.7M_{p}$, while the value of the Higgs field at the beginning of inflation is obtained by evaluating (13) at $\phi_{0}$.
 3 Scalar fields in the Einstein frame
@@
 In other words, we get started in the Jordan frame and after conformal transformation, the matter sector is gravitationally coupled to the inflationary sector.
+This aims to induce direct couplings between the background composed of the fields $\phi,h$ and the scale factor $a(t)$, while also making it easier to solve the system’s dynamics.
 In the Jordan frame our model is given by
@@
 In this work, we will consider a massive spectator scalar field $\chi$ that only interacts gravitationally.
-The action of such a field is given by
-$S_{J}[\chi]=\int\mathrm{d}x^{4}\sqrt{-g_{J}}\left(-\dfrac{1}{2}g^{\mu\nu}_{J}\partial_{\mu}\chi\partial_{\nu}\chi-\dfrac{m_{\chi}^{2}}{2}\chi^{2}\dfrac{1}{2}\,\xi_{\chi}\,R_{J}\,\chi^{2}\right)\,.$ (29)
+To the reader interested in the evolution and production of gauge bosons and other kind of matter fields after inflation, we recommend [37] and the references therein.
+The action in the Jordan frame of a free scalar field $\chi$ is given by
+$S_{J}[\chi]=\int\mathrm{d}x^{4}\sqrt{-g_{J}}\left(-\dfrac{1}{2}g^{\mu\nu}_{J}\partial_{\mu}\chi\partial_{\nu}\chi-\dfrac{m_{\chi}^{2}}{2}\chi^{2}-\dfrac{1}{2}\,\xi_{\chi}\,R_{J}\,\chi^{2}\right)\,.$ (23)
 The dimensionless parameter $\xi_{\chi}$ represents the non-minimal coupling to the gravity sector and may take any value; however, studies based on electroweak vacuum stability [45] suggest that the value of non-minimal coupling constant should not exceed $10$.
 Moreover there are two interesting special cases: $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$.
-The value $\xi_{\chi}=0$ is called the case of minimal coupling.
-while $\xi_{\chi}=1/6$ corresponds to conformal coupling.
+The value $\xi_{\chi}=0$ is called the case of minimal coupling while $\xi_{\chi}=1/6$ corresponds to conformal coupling.
 The first case corresponds to a free scalar field theory and the second one ensures the conformal invariance of the model if, in addition, $m_{\chi}=0$.
@@
 They can explain the particle production process, with the gravitational field responsible for creating $\chi$ particles.
-This process of particle creation due to gravitational interactions has implications for the reheating process [17, 41, 42].
-It also plays a role in gravitational dark matter production [43, 44, 45, 12, 11, 9, 46].
-What follows is to perform the conformal transformation (4) to the action (29) to change to the Einstein frame.
-Since $\chi$ is a spectator field, it remains energetically subdominant, $\rho_{\chi}\ll 3H^{2}M_{p}^{2}$, throughout the entire background evolution.
-Therefore, we can apply the same conformal transformation of the metric (4) as in the previous section.
+This process of particle creation due to gravitational interactions has implications for the reheating process [17, 46, 47], and it also plays a role in gravitational dark matter production [48, 49, 50, 12, 11, 9, 51].
+However, as pointed out in [52], the analysis of matter fields with non-minimal coupling in multi-field models of inflation becomes complicated even if the field metric is flat, as moving from one conformal frame to another induces non-canonical kinetic terms.
+The interpretation of the reheating phase after inflation also changes.
+While in the Jordan framework it is mainly due to the oscillatory behavior of the Hubble parameter, in the Einstein framework it is mainly due to the explicit interaction terms between the matter fields, and the inflationary sector of the model, through which particle production and reheating of the universe occurs.
+Despite these differences, evaluating the inflationary dynamics is much simpler in the Einstein frame [53], as in this frame, Einstein’s equations are second-order, making their evaluation more straightforward [54, 55].
+Since the computation of the $\beta$-Bogolyubov coefficient is independent of the conformal frame [3], we choose to perform the calculation in the Einstein frame to avoid unnecessary complications when studying the background dynamics in the Jordan frame.
+While gravitational particle production is typically carried out in the Jordan frame [56, 57, 9, 58], performing the analysis in the Einstein frame [52, 59, 60] is perfectly valid and yields similar results.
+As we assume that the energy density of the $\chi$ field never dominates, the reheating phase after inflation cannot proceed purely through gravitational interactions.
+Instead, it is assumed that reheating occurs via the decay of the scalaron into light particles $m_{\chi}\ll M$ [61, 8, 62] and through the production of gauge bosons due to Higgs boson decay [28, 37, 35].
+Thus, the spectator field $\chi$ is considered as dark matter produced gravitationally.
+Given that the $\chi$-field interacts only gravitationally with the background fields $\phi$ and $h$, its dynamics can be analyzed effectively within the Einstein frame.
+To study its evolution, we apply the same conformal transformation of the metric (4) as in the previous section.
 By doing so, we avoid considering back-reaction effects on the space-time geometry and the background dynamics.
-In the Einstein frame, we obtain the following action
+In the Einstein frame, we obtain the following action:
 $S_{E}[\tilde{\chi}]=-\dfrac{1}{2}\int\mathrm{d}^{4}x\,\sqrt{-g}\left(g^{\mu\nu}\partial_{\mu}\tilde{\chi}\partial_{\nu}\tilde{\chi}+\xi_{\chi}R\tilde{\chi}^{2}\right.\\[8.5359pt] +\left.e^{-\alpha\phi}m_{\chi}^{2}\tilde{\chi}^{2}\right)+3\alpha\left(\xi_{\chi}-\dfrac{1}{6}\right)\int\mathrm{d}^{4}x\,\sqrt{-g}\left[\tilde{\chi}\,g^{\mu\nu}\partial_{\mu}\tilde{\chi}\partial_{\nu}\phi\right.\\[8.5359pt] \left.+\dfrac{\alpha}{4}\tilde{\chi}^{2}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi\right]\,,$ (24)
@@
 We can perform another conformal transformation which involves the scalar field $\tilde{\chi}$ leading to another Einstein frame to remove the non-minimal coupling constant $\xi_{\chi}$.
-However, we are only interested in the situation where $\chi$ is a spectator field and both scalaron field $\phi$ and Higgs field are responsible for the dynamics during and after inflation.
+However, we are only interested in the situation where $\chi$ is a spectator field and both scalaron field $\phi$ and Higgs field $h$ are responsible for the dynamics during and after inflation.
 One can see that the conformal transformation (4) induces interaction terms between the scalaron $\phi$, the Higgs field $h$, and the free scalar $\tilde{\chi}$.
@@
 This action describes the interaction between the background fields $\phi$ and $h$ with the massive spectator field $\tilde{\chi}$.
-We can see that even with $\xi_{\chi}=0$ this interaction does not disappear, unless the field $\tilde{\chi}$ is conformal invariant ($m_{\chi}=0\,,\xi_{\chi}=1/6$).
-In the paper [8], the authors study the gravitational production of dark matter in the $R^{2}$ inflation model and show that the scalaron $\phi$ can decay perturbatively in dark matter particles $\chi$ with masses $m_{\chi}<M/2$, where $M\simeq 1.3\times 10^{-5}M_{p}$ is the mass term of the scalaron.
-For heavy particles, perturbative particle production is kinematically forbidden.
-However, the production of scalar particles is not limited to perturbative methods; there may be particle production due to the expansion of the universe and the dynamics of the fields involved during the inflationary epoch.
-In our case, the classical time-dependent background is defined by the evolution of the scale factor $a(t)$ and the evolution of the scalar fields $\phi$ and $h$.
-That is, the matter-free scalar field $\chi$ propagates in a time-dependent classical background, whose dynamics are given by the evolution of the $\phi$ and $h$ fields of the Higgs-$R^{2}$ model.
-To study this process it is necessary to find the equation of motion for the spectator field $\tilde{\chi}$ and track the mode evolution during and after inflation.
-Varying the action (30) respect to $\tilde{\chi}$ yields the following equation of motion
+We can see that even with $\xi_{\chi}=0$, this interaction does not disappear unless the field $\tilde{\chi}$ is conformally invariant ($m_{\chi}=0\quad\xi_{\chi}=1/6$).
+Since we are considering $\chi$ as a spectator dark matter field that interacts only gravitationally, it is natural to explore its production mechanisms within inflationary models.
+In the literature, there are studies on the production of dark matter in the $R^{2}$ model [8, 63, 62, 64], where a scalar field $\chi$ with action (23) is considered.
+In [8, 63, 62], the authors study the gravitational production of dark matter in the $R^{2}$ inflation model and show that the scalaron $\phi$ can decay perturbatively into dark matter particles $\chi$ with masses $m_{\chi}\ll M$, where $M\simeq 1.3\times 10^{-5}M_{p}$ is the mass term of the scalaron.
+The corresponding two-body decay width is given by [61, 8, 63].
+$\Gamma_{\phi\to\chi\chi}=\dfrac{1}{192\pi M_{p}^{2}}\left(1-6\xi_{\chi}+2\dfrac{m_{\chi}^{2}}{M^{2}}\right)^{2}\sqrt{1-\dfrac{4m_{\chi}^{2}}{M^{2}}}$ (26)
+One can see that for heavy particles $m_{\chi}\gtrsim M$, perturbative particle production is kinematically forbidden.
+The production of light particles after inflation occurs mainly through this mechanism [8].
+However, in [64] the authors study the gravitational production fo heavy dark matter particles with minimal coupling during and after inflation.
+They find that no additional particle production occurs during inflation if the particle mass is smaller than the Hubble scale $H_{\mathrm{inf}}$ during inflation.
+However, if the mass is larger, the production after inflation becomes dominant.
+On the other hand, there have been few attempts to study spectator dark matter field in the Higgs-$R^{2}$ model.
+In [65], the authors find that the Higgs-$R^{2}$ model is equivalent to a scalar singlet dark matter or Higgs-portal model, constraining the scalar singlet dark matter sector using the dark matter abundance and model parameters.
+This work, however, does not consider an additional scalar field coupled to the background.
+In [66], the authors introduce an extra scalar degree of freedom with non-minimal coupling that plays the role of freeze-in dark matter.
+They also discuss the freeze-in production of dark matter both from the non-thermal scattering during reheating and the thermal scattering after reheating.
+In our case, we investigate the gravitational production of the $\chi$ field through non-perturbative mechanisms, taking into account the induced interaction terms in the Einstein frame.
+Specifically, particle production can occur due to the expansion of the universe and the dynamics of the fields involved during the inflationary epoch.
+The classical time-dependent background in this scenario is determined by the evolution of the scale factor $a(t)$ and the scalar fields $\phi$ and $h$.
+That is, the spectator field $\chi$, which is initially decoupled from other matter fields, propagates in a time-dependent classical background governed by the evolution of $\phi$ and h in the Higgs-$R^{2}$ model.
+To analyze this process, we need to derive the equation of motion for the spectator field $\tilde{\chi}$ and track the evolution of its modes during and after inflation.
+By varying the action (24) with respect to $\tilde{\chi}$, we obtain the following equation of motion:
 $(\square-e^{-\alpha\phi}\,m_{\chi}^{2}-\xi_{\chi}\,R)\tilde{\chi}+\left[\dfrac{\alpha}{2}\,\left(1-6\xi_{\chi}\right)\square\phi-\dfrac{\alpha^{2}}{4}\left(1-6\xi_{\chi}\right)(\partial\phi)^{2}\right]\tilde{\chi}=0\,.$ (27)
@@
 Consequently, in the Einstein frame, the model presents an interaction term with the dynamics of the classical background through the Ricci scalar, the scale factor $a$, and the dynamics of the fields $\phi$ and $h$.
-As a result, studying the solutions is generally highly complex from an analytical perspective; hence a numerical treatment of the background dynamics is necessary for a comprehensive analysis of mode evolution, especially after inflation when both fields $\phi$ and $h$ exhibit a strong interacting anharmonic oscillatory behavior, as one can see in the figure 2.
+As a result, studying the solutions is generally highly complex from an analytical perspective; hence a numerical treatment of the background dynamics is necessary for a comprehensive analysis of mode evolution, especially after inflation when both fields $\phi$ and $h$ exhibit a strong interacting anharmonic oscillatory behavior, as one can see in the Fig.
+2.
 By numerically determining the background evolution of the scalaron, the Higgs field, and the scale factor, we can compute the Ricci scalar using $R=6(\dot{H}+2H^{2})$, ensuring that the dynamics are well-determined.
@@
 where $\Omega^{2}(t)$ is a background variables dependent and is given by
-$\Omega^{2}(t)=e^{-\alpha\phi}m_{\chi}^{2}-a^{-2}\nabla^{2}+\xi_{\chi}R\\[5.69046pt] +3\alpha\left(\dfrac{1}{6}-\xi_{\chi}\right)(\ddot{\phi}+3H\dot{\phi})-\dfrac{3\alpha^{2}}{2}\left(\dfrac{1}{6}-\xi_{\chi}\right)\dot{\phi}^{2}\,.$ (34)
+$\Omega^{2}(t)=e^{-\alpha\phi}m_{\chi}^{2}+\dfrac{k^{2}}{a^{2}}+\xi_{\chi}R+3\alpha\left(\dfrac{1}{6}-\xi_{\chi}\right)(\ddot{\phi}+3H\dot{\phi})-\dfrac{3\alpha^{2}}{2}\left(\dfrac{1}{6}-\xi_{\chi}\right)\dot{\phi}^{2}\,.$ (29)
 Rescaling the field $\tilde{\chi}\to a^{-3/2}(t)\,\tilde{\chi}(t,\mathbf{r})$ and using the background equations of motion (16), (19) the equation for the scalar field $\tilde{\chi}$ becomes a oscillator-like equation with time-dependent frequency
@@
 This occurs when the modes are well inside the horizon at the beginning of inflation.
-In the figure 3 we show the evolution of $\omega_{k}^{2}(t)$ for a non-minimal coupling $\xi=0$.
+In the Fig.
+3 we show the evolution of $\omega_{k}^{2}(t)$ for a non-minimal coupling $\xi=0$.
 Before $t_{\mathrm{cross}}$, the frequency is dominated by the decay of the physical mode $k/a\sim e^{-Ht}$.
@@
 The adiabaticity condition is given by the adiabatic parameter $A_{k}$ defined as
+Figure 4: An example of non-adiabatic mode evolution leading the gravitational particle production of $\chi$ particles described by the adiabaticity parameter $|A_{k}(t)|$ for $m_{\chi}=10^{-2}H_{\mathrm{end}}$ and $\xi_{\chi}=1/6$.
+At the end of inflation, one sees an abrupt increase on $|A_{k}(t)|$ and non-adiabaticity is induced, leading to gravitational particle production $n_{k}(t)\neq 0$.
+At late times, when the stabilization occurs, we recover the adiabaticity evolution.
 $A_{k}(t)=\dfrac{\dot{\omega}_{k}}{\omega_{k}^{2}}\,.$ (39)
@@
 While this condition is met, no particle production occurs, as adiabatic evolution is ensured, and there is no mixing of negative- and positive-frequency modes.
-From figure 3, we can see that the evolution of $\omega_{k}(t)$ is smooth until the end of inflation.
-At that point, the condition $|A_{k}|\ll 1$ is strongly violated, leading to an enhancement in particle production, and gravitational particle production is enhanced at the time $t_{*}$, corresponding to the end of inflation.
-We show the non-adiabaticity evolution in the figure 4 parametrized by the adiabatic parameter $|A_{k}(t)|$.
+From Fig.
+3, we can see that the evolution of $\omega_{k}(t)$ is smooth until the end of inflation.
+At that point, the condition $|A_{k}|\ll 1$ is strongly violated, leading to an enhancement in particle production, and gravitational particle production is enhanced at the time $t_{*}$, corresponding to the end of inflation $t_{\mathrm{end}}$.
+We show the non-adiabaticity evolution in Fig.
+4 parametrized by the adiabatic parameter $|A_{k}(t)|$.
 The mixture of positive and negative frequency modes at time $t_{\mathrm{end}}$ can be represented as the sum of the solutions $\exp(\pm i\int^{t}\omega_{k}(t^{\prime})\mathrm{d}t^{\prime})$ using the WKB approximation [14, 5]
-$X_{k}(t)=\dfrac{\alpha(t)}{\sqrt{2\omega_{k}(t)}}\,e^{-i\int^{t}\omega_{k}(t^{\prime})\mathrm{d}t^{\prime}}+\dfrac{\beta(t)}{\sqrt{2\omega_{k}(t)}}\,e^{i\int^{t}\omega_{k}(t^{\prime})\mathrm{d}t^{\prime}}\,,$
+$X_{k}(t)=\dfrac{\alpha_{k}(t)}{\sqrt{2\omega_{k}(t)}}\,e^{-i\int^{t}\omega_{k}(t^{\prime})\mathrm{d}t^{\prime}}+\dfrac{\beta_{k}(t)}{\sqrt{2\omega_{k}(t)}}\,e^{i\int^{t}\omega_{k}(t^{\prime})\mathrm{d}t^{\prime}}\,,$ (41)
 where $\alpha_{k}(t)$ and $\beta_{k}(t)$ are the time-dependent Bogolyubov coefficients.
 Comparing with the Bunch-Davies initial condition (40), one see thats $\alpha_{k}(t_{0})=1$ and $\beta_{k}(t_{0})=0$.
-Figure 4: An example of non-adiabatic mode evolution leading the gravitational particle production of $\chi$ particles described by the adiabaticity parameter $|A_{k}(t)|$ for $m_{\chi}=10^{-2}H_{\mathrm{end}}$ and $\xi_{\chi}=1/6$.
-At the end of inflation, one sees an abrupt increase on $|A_{k}(t)|$ and non-adiabaticity is induced, leading to gravitational particle production $n_{k}(t)\neq 0$.
-At late times, when the stabilization occurs, we recover the adiabaticity evolution.
-In other words, gravitational particle production occurs at very early times $t_{\mathrm{in}}\to 0$ and at late times $t_{f}\to\infty$ if the initial and final vacuum states are not the same.
+Gravitational particle production is a quantum process that occurs in the early universe, primarily near the end of inflation and continues for a few e-folds afterwards until the $\beta$-Bogolyubov coefficient is stabilized, allowing for a well-defined computation of the energy density of the produced particles.
+However, if the initial and final vacuum states are not the same, particle production can also be interpreted in terms of asymptotic in and out states, highlighting its connection to both early and late-time behavior.
 We define the initial vacuum state with the corresponding creation-annihilation operators $\hat{a}^{\dagger}_{\mathbf{k}}$ and $\hat{a}_{\mathbf{k}}$ by
@@
 $|\beta_{k}(t)|^{2}=\dfrac{\omega_{k}}{2}|X_{k}|^{2}+\dfrac{|\dot{X}_{k}|^{2}}{2\omega_{k}}-\dfrac{1}{2}\,.$ (44)
-The physical number density $n_{\chi}(t)$ in the late-time limit $t\to\infty$ is given by
-$a^{3}(t)\,n_{\chi}(t)=\int_{0}^{\infty}\mathrm{d}\log k\,n_{k}(t)\,,\quad n_{k}(t)=\dfrac{k^{3}|\beta_{k}(t)|^{2}}{2\pi^{2}}\,,$
-where $n_{k}(t)$ is the comoving number density of particles with comoving wavenumber $k$.
+The total number comoving density $n_{\chi}(t)$ of created particles in the late-time limit $t\to\infty$ is given by
+$a^{3}(t)\,n_{\chi}(t)=\int_{0}^{\infty}\dfrac{\mathrm{d}^{3}k}{(2\pi)^{3}}|\beta_{k}(t)|^{2}\,,$ (45)
+which can be rewritten as
+$a^{3}(t)n_{\chi}(t)=\int\dfrac{\mathrm{d}k}{k}n_{k}(t)\,,\quad n_{k}(t)=\dfrac{k^{3}}{2\pi^{2}}|\beta_{k}(t)|^{2}$ (46)
+where $n_{k}(t)$ is the comoving spectral number density of particles.
 Notice that the expression (44) is only valid when $\omega_{k}^{2}(t)>0$, so the comoving number density of particles created is valid at times when the frequency (34) is positive.
+In some cases, it is possible to analytically evaluate the $\beta$-Bogolyubov coefficient.
+The expression (44) is typically used in numerical approaches, as, in most cases, evaluating $\beta_{k}$ numerically requires knowledge of the mode dynamics $X_{k}(t)$ and the frequency $\omega_{k}(t)$.
+That is, we need to numerically solve the equation (33) with Bunch-Davies initial conditions.
+However, there is an alternative method to compute the Bogolyubov coefficients $\alpha_{k}$ and $\beta_{k}$.
+The WKB approximation (41) provides a solution to (33) if the time-dependent Bogolyubov coefficients $\alpha_{k}(t)$ and $\beta_{k}(t)$ satisfy the system of equations given in [14].
+$\dot{\alpha}_{k}=\dfrac{\dot{\omega}_{k}}{2\omega_{k}}e^{2i\psi_{k}(t)}\,\beta_{k}\,,\quad\dot{\beta}_{k}=\dfrac{\dot{\omega}_{k}}{2\omega_{k}}e^{-2i\psi_{k}(t)}\,\alpha_{k}\,,$ (47)
+where $\psi_{k}(t)=\int_{t_{i}}^{t}\mathrm{d}t^{\prime}\,\omega_{k}(t^{\prime})$ is a time-dependent oscillatory phase.
+Gravitational particle production occurs in the Bunch-Davies regime, where $|\beta_{k}|\ll 1$ and $|\alpha_{k}|\simeq 1$.
+In this case, the Bogolyubov coefficient $\beta_{k}$ is given by
+$|\beta_{k}(t)|\simeq\dfrac{1}{2}\int_{t_{i}}^{t}\mathrm{d}t^{\prime}\,\dfrac{\dot{\omega}_{k}}{\omega_{k}}\exp\left(-2i\int_{t_{i}}^{t^{\prime}}\mathrm{d}t^{\prime\prime}\omega_{k}(t^{\prime\prime})\right)\,.$ (48)
+It is possible to numerically solve this integral without the need to numerically solve the equation (33) for an initial time $t_{i}\ll t$; however, the numerical calculation of the integral becomes complex in most cases when the oscillatory phase $\psi_{k}(t)$ oscillates rapidly.
+This approach is more useful when one seeks to evaluate $\beta_{k}$ analytically or approximately.
+To do so, the steepest descent method is commonly employed in the literature [10, 12, 64, 68].
+We have chosen to adopt a numerical approach to compute (45), due to the nature of the model, which involves a non-minimal coupling and a nontrivial potential term.
+These features significantly complicate the evaluation of the integral expression (48).
+As this analysis lies beyond the main scope of the present work, we leave it for future investigations.
 5.1 Numerical results
@@
 For example, for $N_{*}=60$ at the time of crossing, the initial time corresponds to $t_{\mathrm{ini}}\approx 28.712$ in units of $[10^{-5}M_{p}^{-1}]$, and the corresponding mode $k=aH$ will be the value of $aH$ at time $t_{\mathrm{ini}}$.
-Then, the Bogolyubov coefficient $|\beta_{k}(t)|^{2}$ is calculated using (49), and the comoving number density of gravitational particle production is determined by $n_{k}(t)$.
+In this way, we can associate a value of $k$ to each value of $N_{e}$, which determines the scale $k$ that crosses the horizon $N_{e}$ e-folds before the end of inflation.
+Figure 5: The Bogolyubov coefficient $\beta_{k}$ is shown as a function of cosmic time $t$ for a spectator field with masses $m_{\chi}=10^{-1}H_{\mathrm{end}}$ and $m_{\chi}=3H_{\mathrm{end}}$, corresponding to a mode that exits the horizon $N_{k}=60$ e-folds before the end of inflation.
+We observe that $\beta_{k}$ increases near the end of inflation and eventually stabilizes at late times, for $t\gg t_{\mathrm{end}}$.
+An alternative approach to selecting the $k$-modes is to work directly with the dimensionless ratio $k/k_{\mathrm{end}}$ in the equation for $X_{k}(t)$.
+In this framework, modes with $k<k_{\mathrm{end}}$ will have crossed the horizon at some point before the end of inflation.
+However, for our purposes, it is more convenient to generate a list of modes that satisfy $k=aH$ by solving the background equations and associating each mode with the corresponding number of e-folds $N_{e}$ before the end of inflation.
+The resulting list of physical modes $k$ is then normalized by $k_{\mathrm{end}}$ to construct the spectrum $n_{k}(t)$.
+The Bogolyubov coefficient $|\beta_{k}(t)|^{2}$ is computed using Eq.
+(44), and the comoving number density of gravitationally produced particles is determined by $n_{k}(t)$.
+It is important to note that the particle number is evaluated only once it has stabilized; that is, we extract the value of $n_{k}$ at the end of the numerical simulation to ensure that it has reached a steady state.
 5.1.1 Conformal coupling $\xi_{\chi}=1/6$
 We first present the results for conformally coupled particles ($\xi_{\chi}=1/6$).
-As an example, we present the time evolution of the comoving number of particles created $n_{k}(t)$ as a function of time for two different cases: light particles $m_{\chi}<H_{\mathrm{end}}$ and heavy particles $m_{\chi}\gtrsim H_{\mathrm{end}}$.
-We should note that the expression for $\beta_{k}(t)$ is only valid when $\omega_{k}^{2}(t)>0$ , we point out that $n_{k}(t)$ only makes physical sense at late times $t\to\infty$ [5].
+As an example, we present the time evolution of the $\beta$-Bogolyubov coefficient as a function of time for two different cases: light particles $m_{\chi}<H_{\mathrm{end}}$ and heavy particles $m_{\chi}\gtrsim H_{\mathrm{end}}$.
+We should note that the expression for $\beta_{k}(t)$ is only valid when $\omega_{k}^{2}(t)>0$ , we point out that $n_{k}(t)$ only makes physical sense at late times $t\to\infty$ [5] when its value stabilizes.
+Figure 6: Numerical $n_{k}/k_{\mathrm{end}}^{3}$ vs $k/k_{\mathrm{end}}$ spectra of conformally-coupled scalars for differente light masses $m_{\chi}$.
+One can see that the UV tail is suppressed for this range of masses.
 The frequency (34) for scalar conformally coupled turns out to be
@@
 For this reason, we take into account the sub-horizon evolution of the modes $X_{k}(t)$ and we take $k$ deep inside the horizon at the beginning of the simulation.
-In the figure, we present two cases: $m_{\chi}=10^{-5}H_{\mathrm{end}}$ and $m_{\chi}=10^{-1}H_{\mathrm{end}}$.
-In both panels of figure 5, we observe that the number of particles created saturates at the end of inflation, stabilizing and becoming constant at late times $t\to\infty$.
-This growth in $n_{k}$ is due to the non-adiabaticity of $\omega_{k}$ at the end of inflation, as can be seen in figure 4.
-Figure 5: Comoving number density of particles, plotted on a logarithmic scale, as a function of cosmic time $t$ for two different values of the spectator mass $m_{\chi}$, with the same mode $k$, corresponding to $N_{k}=60$.
-The graph depicts the evolution of the occupation number $n_{k}/k_{\mathrm{end}}^{3}$, where $k_{\mathrm{end}}=a_{\mathrm{end}}H_{\mathrm{end}}$ denotes the mode at the end of inflation
-The oscillations of $n_{k}(t)$ after inflation arise from resonances occurring at specific moments between the oscillations of the modes $X_{k}$ and the frequency itself.
-During this epoch, particle production is driven primarily by the oscillatory behavior of the scalaron and the Higgs field, as illustrated in figure 2, through the term $\dot{H}$ and the Ricci scalar $R$ via the non-minimal coupling $\xi_{\chi}=1/6$.
-Additionally, we observe that particle production near the end of inflation is most efficient for light particles.
-In both cases, the particle number is well-defined, grows from the end of inflation, and stabilizes at late times, as expected.
-Figure 6: Numerical $n_{k}/k_{\mathrm{end}}^{3}$ vs $k/k_{\mathrm{end}}$ spectra of conformally-coupled scalars for differente light masses $m_{\chi}$.
-One can see that the UV tail is suppressed for this range of masses.
+We present the two different cases in the Fig.
+5: $m_{\chi}=10^{-1}H_{\mathrm{end}}$ and $m_{\chi}=3H_{\mathrm{end}}$.
+One can observe that the number of particles created saturates at the end of inflation, stabilizing and becoming constant at late times $t\to\infty$.
+This growth in $\beta_{k}(t)$ is due to the non-adiabaticity of $\omega_{k}$ at the end of inflation, as can be seen in Fig.
+4.
+The oscillations of $\beta_{k}(t)$ after inflation arise from resonances occurring at specific moments between the oscillations of the modes $X_{k}$ and the frequency itself.
+During this epoch, particle production is driven primarily by the oscillatory behavior of the scalaron and the Higgs field, as illustrated in Fig.
+2, through the term $\dot{H}$ and the Ricci scalar $R$ via the non-minimal coupling $\xi_{\chi}=1/6$.
+For both cases, the particle number is well-defined, grows from the end of inflation, and stabilizes at late times, as expected.
+We can also observe that particle production is more efficient for lighter species, as indicated by the value of $|\beta_{k}(t)|^{2}$ once particle production has ceased.
+In particular, we find $|\beta_{k}(t_{\mathrm{final}})|^{2}\approx 0.0071$ for the case $m_{\chi}=10^{-1}H_{\mathrm{end}}$, while for $m_{\chi}=3H_{\mathrm{end}}$ we obtain $|\beta_{k}(t_{\mathrm{final}})|^{2}\approx 1.216\times 10^{-6}$.
+This behavior is precisely what one expects: heavier particles are produced less efficiently via gravitational mechanisms, as more energy is required to excite them from the vacuum.
 Figure 7: Right: Comoving number density of particles as a function fo cosmic time $t$.
@@
 One can see that the UV tail is suppressed for this range of masses.
-We have numerically computed the spectra of the comoving number density $n_{k}(t)$ as a function of $k/k_{\mathrm{end}}$ for a range of light masses varying from $10^{-5}H_{\mathrm{end}}$ to $10^{-1}H_{\mathrm{end}}$.
-Our findings are summarized in figure 6.
+We have also numerically computed the spectra of the comoving number density $n_{k}(t)$ as a function of $k/k_{\mathrm{end}}$ for a range of light masses varying from $10^{-5}H_{\mathrm{end}}$ to $10^{-1}H_{\mathrm{end}}$.
+Our findings are summarized in Fig.
+6.
 From the figure, one can observe several remarkable features.
@@
 Therefore, the dynamics of the $X_{k}(t)$ modes are dictated by the effective mass term (36) with a quasi-De Sitter spacetime as background with $H\simeq\mathrm{constant}$.
-It is important to note that the UV modes do not cross the horizon during inflation.
+This results in a comoving number density with a scaling $n_{k}\sim k^{3}$ for these modes.
+That is, in the IR region, the particle occupation number vanishes $n_{k}\to 0$ as $k\to 0$, consistent with what is observed in the figure.
+On the other hand, UV modes correspond to those that never crossed the horizon during inflation, and therefore their production is strongly suppressed.
 In place, these modes are mainly produced during the transition to the matter-dominated era and in general during the reheating era.
-The suppression of UV modes occurs because the $\chi$ scalars become supermassive at the end of inflation due to the exponentially scalaron-dependence mass, causing the effective mass to become substantial.
-As a result, gravitational particle production becomes less efficient, leading to a significant suppression in the particle density number $n_{k}(t)$.
-The mass of the field plays a key role in determining the amplitude of $n_{k}$ at the peak momentum $k_{\mathrm{peak}}$, which systematically decreases as $m_{\chi}$ increases [43].
+This behavior is characteristic of gravitational particle production and has been extensively reported in the literature [69, 70, 49, 48, 50].
+The field’s mass $m_{\chi}$ plays a key role in determining the amplitude of $n_{k}$ at the peak momentum $k_{\mathrm{peak}}$, which systematically decreases as $m_{\chi}$ increases [48, 50, 51].
 For large masses $m_{\chi}\gtrsim H_{\mathrm{end}}$, the situation is quite similar.
 As the mass increases, the particle production tends to decrease due to the increase of the effective mass.
-This can be seen in the left panel of figure 7 for a mass of $m_{\chi}=H_{\mathrm{end}}$ and for a mode crossing the horizon at $N_{k}=60$.
-In the rigt panel, we present the comoving spectra of particle production for $m_{\chi}\gtrsim H_{\mathrm{end}}$.
-We find that the behaviour for UV modes is different than in the case of light masses.
-The location of the peak $k_{\mathrm{peak}}$ is different as a function of the mass value, once $k_{\mathrm{end}}$.
-While $k_{\mathrm{peak}}>k_{\mathrm{end}}$ as in the case of light particles, the entry to the horizon, and hence the tail suppression of the UV modes is different.
-This is largely due to the possible resonances that exist for large masses, as the mass term $e^{-\alpha\phi}$ dominates after inflation when $\phi<M_{p}$, which can excite different modes after inflation.
+This behavior can be clearly seen in the particle spectrum $n_{k}$ shown in Fig.
+7.
 Figure 8: The comoving spectrum of gravitational scalar particle production $n_{k}$ for $\xi_{\chi}=0$ as a function of rescaled horizon modes $k/k_{\mathrm{end}}$.
 One can see that the spectrum is nearly scale-invariant for masses $m_{\chi}\lesssim H_{\mathrm{end}}$ and blue-tilted behavior for larger masses
+The case with conformal coupling $\xi_{\chi}=1/6$ exhibits a similar behavior to the light mass case $m_{\chi}\ll H_{\mathrm{end}}$: in the infrared (IR) region, the spectrum scales as $n_{k}\sim k^{3}$, while in the ultraviolet (UV) region, particle production is exponentially suppressed.
 5.1.2 Minimal coupling $\xi_{\chi}=0$
@@
 In this case, the derivative potential $V_{\phi}$ and the Hubble parameter $H^{2}$ contribute so that the effective mass $m_{\mathrm{end}}$ is negative.
+During inflation where $\phi\simeq 5.7M_{p}$, $\dot{H}\to 0$ and the mass term $e^{-\alpha\phi}m_{\chi}^{2}\ll H$, so it is negligible.
+In these cases, the dominant term of (50) is $H^{2}$ if the modes are super-Horizon $k\ll aH$.
 Thus, once the modes cross the horizon and for masses $m_{\chi}\ll H_{\mathrm{end}}$, this may lead to $\omega_{k}^{2}(t)<0$, resulting in tachyonic instabilities, as mentioned in the discussion following equation (36) [46].
@@
 This growth of the modes $X_{\mathbf{k}}(t)$ outside the horizon is the main driver of the enhanced particle production for minimally coupled scalars compared to conformally coupled scalars.
+Figure 9: The Bogolyubov coefficient $|\beta_{k}|^{2}$ as a function of cosmic time $t$ for a spectator mass $m_{\chi}=3H_{\mathrm{end}}$ and minimal coupling $\xi_{\chi}=0$.
 The spectrum of the produced particles is also affected, and its behavior depends on the mass $m_{\chi}$.
 To obtain the gravitational spectra, we use the same technique to solve the equation numerically: we choose the $k$-mode inside the horizon, approximately 5 e-folds before it crosses the horizon and show the numerical result for $\beta_{k}(t)$ at those times when $\omega_{k}^{2}>0$.
-In the figure 8, we show the spectrum of created particles for minimally coupled masses.
+In Fig.
+8, we show the spectrum of created particles for minimally coupled masses.
 We observe that the spectrum is nearly scale-invariant for IR modes and for masses $m_{\chi}\lesssim H_{\mathrm{end}}$ before reaching the peak at $k_{\mathrm{peak}}$.
 However, as the mass increases, the spectrum tends to be blue-tilted.
-In the UV region, all modes decay either as a power law or exponentially.
-Figure 9: Comoving number density of particles as a function of cosmic time.
-A similar behavior can be observed in Figure 6: the number of created particles saturates at the end of inflation and stabilizes, becoming constant at late times
+In the UV region, all modes decay either as a power law or exponentially [49].
+As the authors mention, the comoving particle number density (45) can have an IR divergence for $m_{\chi}\ll H_{\mathrm{end}}$ that is governed by a cutoff scale $k_{0}$.
+For $\xi_{\chi}>0$, the integral (45) converges and becomes IR insensitive approximately when $\xi_{\chi}\sim 1/6$.
 A clear difference exists between the cases $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$.
@@
 In other words, for the case $\xi_{\chi}=1/6$, the Ricci scalar plays an important role in describing the evolution of the modes $X_{k}(t)$, which translates into a different spectrum of created particles.
-From figure 9 we can see that the particle production at minimal coupling is much larger than the conformally-coupled case.
+In the Fig.
+9 we present the behavior of $|\beta_{k}(t)|^{2}$ as a function of time for a mass $m_{\chi}=3H_{\mathrm{end}}$ and $N_{k}=60$ for the minimum coupling case $\xi_{\chi}=0$.
+We see the same behavior as in the case $\xi_{\chi}=1/6$, a resonant peak near the end of inflation, subsequent oscillations until reaching a stabilized and constant value at $t\gg t_{\mathrm{end}}$.
+The difference lies in the order of magnitude of $|\beta_{k}(t)|^{2}$, and hence of $n_{k}$.
+We can see that the particle production at minimal coupling is much larger than the conformally-coupled case.
 This is mainly due to the behavior of the spectator field $X_{k}(t)$ during inflation.
 As explained above, the tachyonic instabilities in the case of $\xi_{\chi}=0$ cause the modes $X_{k}(t)$ to grow exponentially.
-This growth of the super-horizon modes is primarily responsible for the enhanced gravitational particle production for minimally coupled scalars.
+This growth of the super-horizon modes is primarily responsible for the enhanced gravitational particle production for minimally coupled scalars [49, 5].
+6 Relic abundance
+If gravitationally produced particles are sufficiently long-lived, they could survive throughout the entire thermal history of the universe up to the present-day $t=t_{0}$, thereby becoming viable dark matter candidates.
+These particles might account for only a fraction of the observed dark matter, or they could potentially explain its entirety, as quantified by the parameter $\Omega_{\mathrm{DM}}$.
+For this reason, it is worthwhile to compute both the total number density $n_{\chi}$ and the present-day abundance $\Omega_{\chi}h^{2}$.
+In this section, we compute the present-day relic abundance of $\chi$ particles produced gravitationally in the Higgs-$R^{2}$ model, summarize our numerical results, and impose the corresponding relic abundance constraints.
+The total number of gravitationally produced particles is characterized by two key parameters, depending on the application: either through the comoving particle number density, defined in Eq.
+(45), or via the present-day density parameter $\Omega_{\chi}$.
+The present-day particle relic abundance is given by
+$\Omega_{\chi}h^{2}=\dfrac{\rho_{\chi}(t_{0})}{\rho_{c}}h^{2}\,.$ (51)
+Figure 10: Total comoving number density $a^{3}n_{\chi}$ in units of $k_{\mathrm{end}}$ as a function of the spectator mass $m_{\chi}/H_{\text{end}}$.
+The left panel shows the result for the conformal coupling case $\xi_{\chi}=1/6$, while the right panel shows the result for minimal coupling $\xi_{\chi}=0$.
+Here, $\rho_{\chi}$ is the total energy density of the produced particles, and $\rho_{c}=3M_{p}^{2}H_{0}^{2}$ is the critical density of the universe today.
+Once the asymptotic value of $n_{k}(t)$ is known, we can compute $\Omega_{\chi}h^{2}$ at the present-day.
+The renormalized total energy density $\rho_{\chi}$ is defined as [5, 9]:
+$\rho_{\chi}(t)a^{4}(t)=\int\dfrac{\mathrm{d}^{3}k}{(2\pi)^{3}}\,\omega_{k}(t)\,|\beta_{k}(t)|^{2}=\int_{0}^{\infty}\dfrac{\mathrm{d}k}{k}\,\omega_{k}(t)n_{k}(t)\,.$ (52)
+For non-relativistic modes $k\ll a\,m_{\chi}$, which are expected to carry most of the energy at late times, we have:
+$\rho_{\chi}(t)a^{4}(t)\simeq\dfrac{m_{\chi}n_{\chi}}{\rho_{c}}\,.$ (53)
+That is, the present-day abundance at $t=t_{0}$ can be computed from:
+$\Omega_{\chi}h^{2}\simeq\dfrac{m_{\chi}n_{\chi}(t_{0})}{\rho_{c}}h^{2}\,.$ (54)
+It is important to emphasize that the mass term $m_{\chi}$ appearing in the frequency (34) is accompanied by an exponential factor $e^{-\alpha\phi}$, and thus it must be evaluated once $\phi$ has stabilized after reheating.
+To perform this calculation, one needs to know the reheating mechanism, the reheating temperature $T_{\mathrm{RH}}$, and its duration.
+However, our particle production calculation is independent of the reheating process, since we assume that the scalar field is a spectator.
+Reheating should take into account the production of gauge bosons $W$ and $Z$ due to Higgs decay and possible resonances.
+First, we show the total comoving number density $a^{3}n_{\chi}$ as a function of $m_{\chi}$ for both $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$ in Fig.
+10.
+In both cases, we observe a decrease in the number of produced particles as the mass increases, as expected.
+The superheavy case $m_{\chi}\gg H_{\text{end}}$ is particularly interesting for minimal coupling.
+In the right panel of Fig.
+10, we observe a significant enhancement in the comoving number density $a^{3}n_{\chi}$ over the explored mass range, reaching a maximum of order $10^{5}$ for $m_{\chi}/H_{\text{end}}\sim 2.5$, which corresponds to $m_{\chi}\simeq 1.5\times 10^{13}\,\text{GeV}$.
+Although this mass is not strictly in the $m_{\chi}\gg H_{\text{end}}$ regime, as the mass increases further, $a^{3}n_{\chi}$ is expected to decay.
+However, for a more detailed study of the $\xi_{\chi}=0$ case, it is necessary to extend the mass range to the superheavy case, which is computationally more demanding, and requires a much finer numerical analysis.
+This is left for future work.
+The conformally coupled case $\xi_{\chi}=1/6$ follows the expected behavior: it exhibits a peak around $m_{\chi}\sim H_{\text{end}}$, and for heavier spectator masses $m_{\chi}\gtrsim H_{\text{end}}$, gravitational production becomes suppressed.
+To calculate the $\Omega_{\chi}$ abundance, it is necessary to understand the mechanism of the reheating in the model.
+As mentioned, our work is independent of this process.
+There are several attempts to explain ths epoch in the Higgs-$R^{2}$ model.
+For example, in [37], the non-perturbative production of $W$ bosons and Higgs boson self-production after inflation is analyzed.
+This allows for the computation of relevant observables such as the reheating temperature $T_{\mathrm{rh}}$.
+The authors conclude that the dominant post-inflationary dynamics of the model is driven by tachyonic instabilities in the effective masses of the gauge bosons, and to a lesser extent, by the tachyonic mass of the Higgs boson itself.
+This tachyonic regime leads to a cosmologically instantaneous preheating, which allows the post-inflationary evolution of the model to be fully determined.
+In light of these results, we can provide an illustrative analytical estimate for the dark matter abundance $\Omega_{\chi}h^{2}$ of the spectator field $\chi$, based on the comoving number density $n_{\chi}$ and the reheating temperature $T_{\text{RH}}$.
+We evaluate the total comoving number density of produced particles at late times, taken as the final time $t_{\text{final}}$ of our numerical simulations, where the factor $a^{3}n_{\chi}$ has reached a constant value.
+Using the asymptotic value of $n_{\chi}$, we can calculate the relic abundance $\Omega_{\chi}h^{2}$ using the amount of expansion from the end of inflation to today.
+To do this, we must rescale the scale factor from the end of inflation to the present-day, taking into account the reheating epoch.
+Taking into account that after inflation, the production of gauge bosons begins due to the decay of the Higgs field [37], the expression (54) can be rewritten as
+$\Omega_{\chi}h^{2}\simeq\dfrac{m_{\chi}n_{\chi}(t_{0})}{\rho_{c}}=\dfrac{m_{\chi}}{3M_{p}^{2}H_{0}^{2}}\dfrac{a^{3}n_{\chi}}{k_{\text{end}}^{3}}\dfrac{k_{\text{end}}^{3}}{a_{0}^{3}}\,,$ (55)
+where $n_{\chi}(t_{0})=a^{3}n_{\chi}/k_{\text{end}}^{3}=H_{\text{end}}^{3}[a^{3}n_{\chi}/k_{\text{end}}^{3}](a_{\text{end}}/a_{0})^{3}$ is the comoving number density at the present-day.
+The factor $[a^{3}n_{\chi}/k_{\text{end}}^{3}]$ comes from calculating (45) numerically, see Fig.
+10.
+The ratio of the scale factors $(a_{\text{end}}/a_{0})^{3}$ corresponds to the amount of expansion from the end of inflation to the present-day.
+Taking into account that after reheating the entropy density $s$ is conserved, the amount of expansion from the end of inflation to the present time can be rewritten as [51, 11]
+Figure 11: Dimensionless relic abundance (54) of $\chi$ particles for the case $\xi_{\chi}=1/6$ as a function of the mass, assuming a reheating temperature $T_{\text{RH}}\sim 10^{9}\,\text{GeV}$.
+The horizontal line indicates the present-day value of the dark matter abundance $\Omega_{\text{CDM}}h^{2}\sim 0.12$.
+$\left(\dfrac{a_{\text{end}}}{a_{0}}\right)^{3}=\dfrac{\pi^{2}g_{*}T_{0}^{3}}{90M_{p}^{2}H_{\text{end}}^{2}}T_{\text{RH}}\,.$ (56)
+Therefore, the abundance (54) of the created particles rescaled to the present-day becomes
+$\Omega_{\chi}h^{2}=\left(\dfrac{\pi^{2}g_{*}T_{0}^{3}}{270M_{p}^{4}H_{0}^{2}}\right)\left(\dfrac{m_{\chi}}{H_{\text{end}}}\right)\left(\dfrac{a^{3}n_{\chi}}{k_{\text{end}}^{3}}\right)H_{\text{end}}^{2}\,T_{\text{RH}}\,,$ (57)
+where $g_{*}\approx 3.91$, $T_{0}=0.234\times 10^{-12}\,\text{GeV}$ and $H_{0}=2.133\times 10^{-42}\,\text{GeV}$ are the numerical values in GeV of the temperature and the Hubble constant at the present-day.
+Thus, the numerical factor of $\Omega_{\chi}h^{2}$ is
+$\dfrac{\pi^{2}g_{*}T_{0}^{3}}{270M_{p}^{4}H_{0}^{2}}\approx 0.1154\times 10^{-28}\,\text{GeV}^{-3}$ (58)
+The expression (57) allows us to calculate the relic abundance from the numerical calculation of the comoving number density $a^{3}n_{\chi}$, and as a function of the mass $m_{\chi}$ and the reheating temperature $T_{\text{RH}}$.
+For example, by fixing $T_{\mathrm{RH}}$, one can estimate the dependence of $\Omega_{\chi}$ on the mass $m_{\chi}$; alternatively, by fixing the ratio $m_{\chi}/H_{\text{end}}$, the abundance becomes proportional to $a^{3}n_{\chi}/k_{\mathrm{end}}^{3}$, isolating the effect of particle production.
+Since the reheating temperature is a fixed prediction of this model, we focus on exploring the dependence of $\Omega_{\chi}$ on the mass $m_{\chi}$.
+In Fig.
+11, we present the relic abundance of $\chi$ particles at the present-day as a function of their mass for $\xi_{\chi}=1/6$, using Eq.
+(57) and $T_{\text{RH}}\sim 10^{9}$ GeV.
+From the figure we find that the maximum allowed mass for the spectator field to account for the observed value $\Omega_{\text{dm}}h^{2}$ without overproduction is $m_{\chi}\simeq 10^{12}\,\text{GeV}$, consistent with other results in the literature for the conformally coupled case [8, 13], although the exact upper bound may vary depending on the specific production mechanism considered.
+We also observe that for light particles $m_{\chi}\ll H_{\text{end}}$, the abundance decreases.
+This is because in this regime the gravitational production is inefficient due to the conformal coupling, which suppresses the amplification of modes during the expansion of the Universe.
+Although these masses fail to explain all of the observed dark matter, they may contribute a subdominant fraction.
+Therefore, the allowed masses are in the range $m_{\chi}\lesssim 10^{12}\,\text{GeV}$ for the conformal case, considering only non-perturbative dark matter production and heavy particles $m_{\chi}\gtrsim H_{\text{end}}$.
 7 Discussion and remarks
-In this study, we investigated the gravitational and non-perturbative production of free scalar particles, $\chi$, in Higgs-$R^{2}$ inflation.
+In this paper, we have investigated the gravitational and non-perturbative production of free scalar particles, $\chi$, in Higgs-$R^{2}$ inflation.
 We started with a free spectator field, $\chi$, interacting only through the gravitational term $\xi_{\chi}R$ in the Jordan frame.
@@
 For light masses and $\xi_{\chi}=0$, we found that the comoving number of particles created was significantly higher compared to the case of conformally coupled particles, primarily due to the tachyonic growth of the modes $X_{k}(t)$.
-We did not attempt a detailed exploration of the parameter space of the Higgs-$R^{2}$ inflation model or the parameters of the scalar field $\chi$.
-Instead, we focused on the Higgs-$R^{2}$ inflation model, whose observables are consistent with Planck data, and considered the possibility of non-perturbative particle production through the direct couplings that arise between the $\chi$ field and the inflationary sector when changing the conformal frame.
-We found that inflation driven by the Higgs field and the scalaron can produce scalar particles in the cases $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$ for a specific range of masses.
-This field could potentially contribute to the abundance of dark matter, though its present-day abundance would need to be calculated.
+While we did not conduct a comprehensive scan of the parameter space for Higgs-$R^{2}$ inflation or the $\chi$ field, we focused on a benchmark scenario consistent with current CMB observations.
+We found that non-perturbative production of spectator scalars with conformal coupling can yield viable dark matter candidates for masses up to $m_{\chi}\sim 10^{12}\,\mathrm{GeV}$.
+Lighter masses fail to account for the full relic abundance but may still contribute a subdominant component.
+The superheavy regime $m_{\chi}\gg H_{\text{end}}$ is particularly intriguing, involving a different production mechanism driven by adiabaticity violation, but is left for future work due to its computational demands.
+Our results underscore the importance of gravitational production mechanisms in extended inflationary scenarios and highlight the viability of free spectator scalars as dark matter candidates, even in the absence of direct couplings to the Standard Model.
 There remain several directions for future research, including the production of spin-$1/2$ and spin-$1$ dark matter, the production of gauge bosons $W^{\pm}_{\mu}$, $Z_{\mu}$, and the implications of particle production on primordial non-Gaussianities in the CMB.
-It would also be interesting to study the isocurvature constraints for light scalar field masses in this model.
+It would also be interesting to study the isocurvature dark matter constraints for light scalar field masses in this model.
 We leave these intriguing topics for future investigation.
@@
 We are very grateful to Dr.
-Marcos Garcia for his insightful comments and input, which made it possible to write the code for this article.
-We would also like to acknowledge Dr.
-Andrew Long for providing an example of numerical code.
-The work of F.P.
-has been sponsored by CONACHyT-Mexico through a doctoral scholarship.
-This work was also partially supported by CONACHyT- Mexico through the support CBF2023-2024-193
+Marcos García for his insightful comments and guidance, which were instrumental in the development of the numerical code used in this work.
+We also thank Dr.
+Andrew Long for generously providing an example of numerical code that served as a valuable reference during the initial stages of this project.
+We would like to express our sincere appreciation to the anonymous referee for their constructive suggestions and careful reading of the manuscript, which helped us significantly improve the clarity and quality of this work.
+The work of F.P. has been supported by CONACyT-México through the doctoral scholarship 803062. This research was also partially supported by CONACyT-México through the grant CBF2023-2024-193.